### v1

# X-ray panorama of the SS 433/W50 complex by SRG/eROSITA

X-ray panorama of the SS 433/W50 complex by SRG/eROSITA
Galactic microquasar SS 433 and the radio nebula W50 surrounding it present a prototypical example of a hyper-Eddington binary system shaping its ambient interstellar medium via energetic outflows. In this paper, we present X-ray observations of the SS 433/W50 complex by the eROSITA telescope onboard the SRG space observatory. These data provide images of the entire nebula characterized by a very large dynamic range and allow spectral analysis of the diffuse X-ray emission. In particular, these data illustrate a close connection between the thermal and non-thermal components of W50 on scales ranging from sub-parsecs, represented by narrow X-ray bright filaments, to the entire extent of the nebula. These data also allow us to fully characterize a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts, but are prominent in very high energy gamma-ray emission. The resulting multifaceted picture of the interaction between energetic outflows and the surrounding medium paves the way for future focused multiwavelength observations and dedicated numerical simulations.
Key Words.:
ISM: jets and outflows – X-rays: binaries – plasmas – acceleration of particles1 Introduction
SS 433 is a unique close stellar binary system in our Galaxy (located at and kpc) composed of a relativistic compact object (a black hole or a neutron star) which accretes matter from its companion at sustainably hyper-Eddington rate (see Fabrika, 2004; Cherepashchuk et al., 2025, for reviews). Having been first listed in the catalog of point-like H emitters (Stephenson & Sanduleak, 1977), it quickly became one of the most studied astrophysical objects thanks to its very peculiar optical spectrum (Margon et al., 1979) featuring pairs of narrow (with half-opening angle ) blue- and red-shifted lines (with respect to their rest frame positions) with the shifts varying periodically in time (see Margon, 1984, for a review of early observational results). A kinematic model associates this emission with a pair of narrow relativistic jets having constant bulk velocity of a quarter of the speed of light and changing direction in a regular precession manner with a period days (Abell & Margon, 1979; Fabian & Rees, 1979). This period is much longer than the period of orbital modulation ( days) and is interpreted as due to precession of a thick accretion disk with the amplitude of 21∘ and the mean axis inclined by the angle to the line of sight (e.g., Katz, 1980; Sarazin et al., 1980; Whitmire & Matese, 1980). Subsequent detection of extended and variable radio emission (Spencer, 1979) perfectly described by the precessing jets pattern fully confirmed the picture of twin relativistic outflows (Hjellming & Johnston, 1981a, b; Blundell & Bowler, 2004). Moreover, the proper motion of the individual radio-emitting blobs allowed one to constrain the distance to the source with a remarkable 10% accuracy, being kpc (Hjellming & Johnston, 1981a; Blundell & Bowler, 2004; Marshall et al., 2013).
Despite indications of both quasi-regular (nutation, Katz et al., 1982) and sporadic (jittering, Kubota et al., 2010) deviations from the predictions of this simple model as well as presence of (relatively short) periods of time when the Doppler shifts of optical emission lines change significantly (e.g., Blundell et al., 2011; Medvedev et al., 2022), all the derived parameters of the model stay remarkably stable over more than 40 years of continuous monitoring observations (Cherepashchuk et al., 2018). This stability likely reflects extreme robustness of the mass transfer occurring in the hyper-Eddington regime (van den Heuvel, 1981; van den Heuvel et al., 2017) with the estimated accretion rate (e.g., Shklovskii, 1981; Cherepashchuk, 1981; Leibowitz, 1984; Fuchs et al., 2006), where is the mass of the compact object and is the Solar mass and is the critical Eddington mass accretion rate. The compact object is most likely a black hole with (Cherepashchuk et al., 2023), although the possibility of it being a neutron star () has also been discussed (e.g. Goranskij, 2011; Medvedev et al., 2018).
Since X-ray emission coming from the innermost parts of the hyper-Eddington accretion disk is expected to be strongly beamed in the axial direction (Shakura & Sunyaev, 1973; Begelman et al., 2006; Poutanen et al., 2007), we are not able to observe it directly, given the high inclination angle of the system. Indeed, although X-ray emission from this source has also been detected early on (Seward et al., 1976; Forman et al., 1978), its luminosity is relatively low compared to the Eddington luminosity (Grindlay et al., 1984). As a result, SS 433 has long been considered as a beamed away prototype of Ultra Luminous X-ray sources (ULXs, e.g., Kaaret et al., 2017, for a review) observed in normal star-forming galaxies (Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007), although possible indications of the ‘hidden’ central emission are yet to be confirmed (Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016, 2019; Middleton et al., 2021; Fogantini et al., 2023).
Instead, what dominates X-ray emission of the system is again thermal emission from a twin pair of the baryonic jets having almost identical orientation and bulk velocity to the optically emitting jets (Watson et al., 1986). Numerous pairs of blue- and red-shifted narrow emission lines of highly ionized heavy elements (primarily, silicon, sulfur, iron and nickel) allow one to build a multi-temperature model of the ballistically expanding and cooling flow (Watson et al., 1986; Kotani et al., 1996; Marshall et al., 2002; Brinkmann et al., 2005; Medvedev & Fabrika, 2010; Marshall et al., 2013; Khabibullin et al., 2016) which first becomes visible when its temperature is above 20 keV and then cools down below 1 keV and likely fragments later on due to thermal instability (Brinkmann et al., 1988). Detailed spectroscopic modeling of this emission (e.g. Khabibullin et al., 2016; Medvedev et al., 2019) in combination with the eclipsing observations of the jets by the companion star (Stewart et al., 1987; Kawai et al., 1989; Filippova et al., 2006; Lopez et al., 2006; Atapin & Fabrika, 2016) enable very robust determination of the jets’ physical properties (so-called jets ‘biometry’, consisting of their size, opening angle, velocity, temperature, and the mass flux). Thanks to this we know very robustly the kinetic luminosity of these outflows, which exceeds (e.g., Medvedev et al., 2019), confirming the necessity of a hyper-Eddington ‘engine’ releasing large amount of energy in the form of relativistic outflows, consistently with the early theoretical (Shakura & Sunyaev, 1973) and modern numerical (e.g., Ohsuga & Mineshige, 2011; Jiang et al., 2019; Yoshioka et al., 2022; Fragile et al., 2025) predictions.
In contrast to the radiation freely escaping the system, the kinetic energy of the launched outflows cannot be easily transported away from the source and has to be deposited into the surrounding interstellar medium. The estimated total amount of this energy exceeds erg for a life-time of the system at the level of 30,000 yr (e.g., van den Heuvel et al., 1980). This energy is at least on par with the energy of the initial supernova explosion that should have accompanied the birth of the compact object, meaning that a quite large region pc in its vicinity can be shaped by this later activity (e.g., Begelman et al., 1980; Zealey et al., 1980; van den Heuvel, 1981; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Churazov et al., 2020, 2024).
Indeed, SS 433 is known to be surrounded by the famous giant ( pc) radio nebula W50 (Westerhout, 1958; Holden & Caswell, 1969; Geldzahler et al., 1980; Elston & Baum, 1987; Dubner et al., 1998; Goodall et al., 2011b; Broderick et al., 2018; Sakemi et al., 2021), which has a suggestive elongated morphology, with the elongation axis being co-aligned with the symmetry axis of the jets precession cone. The asymmetry of the lobes, or “ears”, of the nebula can be readily explained by the noticeable stratification of the surrounding medium, given that the system is located at the height pc below the Galactic plane and has the size comparable to the scale height of the exponential cold gas disk (e.g., Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a). The interior of the nebula is also known to shine in soft thermal X-rays (Brinkmann et al., 1996, 2007; Safi-Harb & Ögelman, 1997; Chi et al., 2024), hard non-thermal (Seward et al., 1980; Watson et al., 1983; Yamauchi et al., 1994; Safi-Harb & Petre, 1999; Namiki et al., 2000; Safi-Harb et al., 2022; Kaaret et al., 2024) and filamentary H emission (Kirshner & Chevalier, 1980; Shuder et al., 1980; Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017; Rosado et al., 2021), indicative of the high velocity shocks propagating through a relatively dense interstellar medium.
The nature of the extended X-ray jets (EXJ) remains more puzzling, although their morphology, spectra, and polarization all point to the synchrotron origin of the emission (e.g., Safi-Harb et al., 2022; Kaaret et al., 2024). Most recently, Very High Energy (VHE) emission from these structures has been detected and mapped (e.g., Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024; LHAASO Collaboration, 2024), demonstrating that indeed ultra-relativistic electrons with energy above 100 TeV are accelerated in the system (e.g, Kimura et al., 2020), probably as a consequence of internal shocks and recollimation in the energetic trans-relativistic outflows (Bykov et al., 2025). The overall production of cosmic rays by such systems (super-Eddington accretors with powerful outflows) might be of importance for the total CR budget and local variations in its spectrum from place to place (e.g., Heinz & Sunyaev, 2002), in particular in the PeV regime (e.g., Peretti et al., 2025; Wang et al., 2025; Bykov et al., 2025).
Although the paradigm of “a super-Eddington X-ray binary powering a nebula with energetic outflows” appears to be well established and consistent with the most salient observational properties of the SS 433/W50 complex, the details of almost all aspects of this interaction remain unclear. Indeed, all the phenomena directly associated with the central source can be trace only up to a distance of mere 0.1 pc from it, where the last signatures of the corkscrew pattern of the precessing jets are observed in radio and X-rays (Migliari et al., 2005; Miller-Jones et al., 2008; Khabibullin & Sazonov, 2017; Sakai et al., 2025). A model invoking steady deceleration and disruption of the jets can, in principle, account for such behavior (Panferov, 2014, 2017), but then the re-appearance of the “jets” with similar to the injection velocities tens of pc away from the central source might be problematic. Morphology of the soft X-ray emission, best mapped by the ROSAT mosaic observations (Brinkmann et al., 1996, 2005), is also quite puzzling, with large global brightness variations and thin filamentary structures, not very commonly observed in similar situations of supernova remnants (which typically feature shell-, plerion- or mixed types of morphology). Deep pointed observations with XMM-Newton, Chandra, Suzaku, and NuSTAR were capable of covering certain parts of the nebula (e.g., Safi-Harb et al., 2022; Chi et al., 2024; Tsuji et al., 2025), but a uniform and sensitive coverage of the entire nebula with adequate spectroscopic capabilities is missing.
Here we present results of X-ray observations of the SS 433/W50 complex by the eROSITA telescope (Predehl et al., 2021) onboard the SRG space observatory (Sunyaev et al., 2021), which were performed during its Performance Verification (PV) phase. The obtained mosaic provides us with X-ray images of the entire extent of W50 with unprecedented dynamic range, bringing X-ray data on par with the radio and other multi-wavelength data, and, in addition, delivers exquisite spectral information for the rich variety of diffuse structures. These data, in particular, demonstrate a close connection between the thermal and non-thermal components of W50 on scales ranging from subparsec, represented by narrow X-ray and radio bright filaments, to the entire extent of the nebula. These data also enable full characterization of a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts.
The structure of the paper is following: we describe the observations and the obtained data in Sect. 2; results of the broad-band and spectrally-resolved imaging are presented in Sect. 3 and Sect. 4, respectively; spectroscopic analysis of the emission from several selected regions follows in Sect. 5; the obtained global picture is discussed in Sect. 6 and summarized in Sect. 7.
2 Data
Since W50 spans approximately degrees on the sky, corresponding to the physical size of at the distance of 5 kpc111The most recent parallax measurement for SS 433 in Gaia DR3 corresponds to the significantly larger distance of kpc (Gaia DR3 4293406612283985024, Gaia Collaboration et al., 2023), we caution of using this value given large residuals of the astrometric solution. The previous parallax measurement published in Gaia DR2 corresponded to a much smaller distance of kpc (Gaia DR2 4293406612283985024, Arnason et al., 2021)., the observing campaign consisted of 9 individual telescope pointings (each covering a circular area of in diameter). These pointings were arranged in a mosaic to result in a nearly uniform coverage of the entire nebula and, in addition, the highest sensitivity in the most interesting regions. The observing dates, from October 5th to October 8th and then from October 19th to October 20th, 2019, were chosen to coincide with the two consecutive orbital eclipses of the central source SS 433 by its normal companion star (according to orbital ephemerids by Eikenberry et al., 2001). During eclipses, the apparent X-ray flux of SS 433 goes down by a factor of several (e.g., Kawai et al., 1989; Lopez et al., 2006; Medvedev et al., 2019), minimizing possible adverse effects of contamination and pile-up due to the presence of a very bright object in the field-of-view, even though relative suppression of the soft X-ray lines being rather minor (Marshall et al., 2013). The typical effective (i.e., equivalent to the nominal eROSITA on-axis performance) exposure time across the map amounts to 30-50 ks, leading to a factor of deeper images compared to the previously available full-extent maps of W50 by observatory in the 0.5-2 keV band (Brinkmann et al., 1996) and Einstein observatory in the 2-4 keV band (Watson et al., 1983). In addition to that, the eROSITA data have much better spatial ( arcsec on the telescope axis and averaged across the field-of-view) and spectral ( eV at 1 keV) resolution (Predehl et al., 2021). A comparable quality of the data is provided by multiple Chandra and XMM-Newton observations, but their coverage of the whole W50 nebula is non-uniform and still incomplete after more than 20 years of observations (Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024). On the other hand, the SRG/eROSITA data are complementary to these observations, as well as to the imaging observations in harder X-rays with NuSTAR, which offer a unique view of non-thermally emitting populations. In this paper, we focus on the SRG/eROSITA X-ray data alone and leave combined analyses for future studies.
We also supplement the PV data with the data obtained in the course of the SRG/eROSITA All-Sky Survey (Predehl et al., 2021; Sunyaev et al., 2021). Although the survey data are much shallower in sensitivity (the typical exposure time per point is s after combining the data of four scans), they provide uniform coverage of the entire area surrounding the nebula. For them, the standard data reduction, background estimation and exposure- and vignetting corrections were applied in the same way as was done for mapping diffuse emission from nearby galaxy clusters (Churazov et al., 2021a, 2023), supernova remnants (Churazov et al., 2021b; Khabibullin et al., 2023, 2024), and the Galactic diffuse emission (Predehl et al., 2020; Khabibullin et al., 2022; Churazov et al., 2024) on even larger scales. Here we also go one step further and combine the PV and all-sky data in a seamless fashion, similarly as it has been done for Chandra and SRG/eROSITA observations of the Perseus cluster (Churazov et al., 2025).
A comparison of the degrees field in radio and X-ray bands is shown in Fig. 1. Although in the radio image, the outer boundary of the W50 nebula appears to be very prominent, the X-ray image is instead dominated by the very bright central object (SS 433) and elongated structures known as “extended X-ray jets” (EXJ) located well inside the radio boundary of W50 (Seward et al., 1980; Watson et al., 1983; Brinkmann et al., 1996). At a distance of 5 kpc, these images span 350 pc and range from 17 pc to almost 370 pc in distance from the Galactic plane. Given that the scale heights for the cold and warm components of the interstellar medium are tens and hundreds of pc, respectively, one could naturally expect signatures of the environment stratification (e.g., Goodall et al., 2011a). This is indeed revealed by the asymmetry of the W50 lobes, with the one directed towards the Galactic plane (the Western Lobe in equatorial coordinates) being more compact than the opposite one (the Eastern lobe).
On larger scales, clear variations of the background and foreground diffuse X-ray emission are seen. Inspection of 3D reddening maps (e.g., Green et al., 2019) shows a clear anti-correlation of the reddening and soft X-ray surface brightness (see Appendix C), suggesting that absorption of X-rays is the main reason for these variations. Morphologically similar structures appear in the 3D reddening maps at a distance of 2-3 kpc, i.e., in the foreground to W50/SS 433. Therefore, it is plausible that the majority of them are physically unrelated to W50.
3 Broad-band imaging
We now turn to the deeper PV data. The broad-band (0.5-4 keV) vignetting-corrected X-ray image of the entire W50 nebula is shown in Fig. 2, with the contours of radio emission at 4.85 GHz (Gregory et al., 1996) overlaid in white. The color coding reflects the measured X-ray surface brightness on a square-root scale spanning two orders of magnitude, highlighting the wide dynamic range of the diffuse X-ray emission captured by eROSITA. In addition to SS 433 itself (the extremely bright source in the center) and a multitude of point sources, the majority of which are nearby foreground stars or background AGN, the picture shows bright diffuse emission with rich internal structure.
The X-ray picture can be broadly decomposed into three components. Firstly, there is a low surface brightness emission filling the entire projected extent of the nebula. It is brighter in the Eastern and Western lobes of W50 (left and right, respectively, in Fig. 2) and significantly dimmer along the Northern and Southern radio boundaries of W50. Remarkably, the central part of W50 (beyond the region contaminated by the PSF wings from SS 433) appears even fainter in X-rays, suggesting a shell-like geometry of the X-ray-emitting gas.
Secondly, one can notice the presence of very narrow X-ray filaments, particularly inside and close to the boundaries of the Eastern and Western lobes. Their apparent width is , while their length is . Given the angular resolution of eROSITA (better than , i.e., at 5 kpc), these filaments are fully resolved in both directions. The X-ray filaments form a complicated network with an overall structure very similar to the filaments of the cold and dense gas observed via H+[N II] emission at different locations within the nebula (Boumis et al., 2007; Farnes et al., 2017). The largest X-ray filaments have a flat transverse profile of the surface brightness and are an order of magnitude brighter than the surrounding medium. Such filaments of the dense X-ray emitting gas might arise as instabilities in the sheared flows, caustics of the strong shock waves, or due to heating of the preexisting cold gas filaments by hot gas and relativistic particles inside the lobes.
The final and perhaps the most visual feature of the image is a pair of nearly symmetrical, sharp-edged “extended X-ray jets” directed away from the central source. Their axis is not only aligned with the elongation of the radio nebula and location of the X-ray bright lobes, but also coincides with the projection of the precession axis of SS 433’s famous trans-relativistic baryonic jets. The physical nature of these structures, first found by the Einstein observatory (Watson et al., 1983) and then observed with all major X-ray telescopes (Brinkmann et al., 1996; Safi-Harb & Ögelman, 1997; Namiki et al., 2000; Brinkmann et al., 2007; Safi-Harb & Petre, 1999), including the most recent polarimteric observations with IXPE (Kaaret et al., 2024), remains a matter of debate (e.g. Churazov et al., 2024). The SRG/eROSITA data clearly show a very rapid onset of emission at the innermost boundary of EXJs on both sides of the nebula (also demonstrated recently by Tsuji et al., 2025, using Chandra data).
4 Spectrally-decomposed imaging
We take advantage of the unique possibility provided by /eROSITA to probe the spectral characteristics of diffuse X-ray emission throughout the full extent of the nebula for the first time. To produce a spectrally-decomposed image, we divided the full spectral band presented earlier into three sub-bands: soft (0.5-1 keV), medium (1-2 keV), and hard (2-4 keV) bands, respectively. A composite RGB image showing the surface brightness of the emission (now on a linear scale) in these bands in red, green, and blue, respectively, is presented in Figure 3. The chosen spectral decomposition immediately reveals the presence of two distinct components in the observed diffuse emission: large regions of red-orange color where emission below 1 keV dominates, and green-cyan structures where prominent emission above 1 keV is present (including the bright central source SS 433). Compact sources also form two groups: the softer (“red-yellow”) ones are predominantly foreground stars, while harder (“blueish”) sources are mostly background AGNs, which shine through the gas and dust in the disk of the Milky Way. Zoom-in images of the lobes (rotated and put side by side to help the comparison) are shown in Fig. 4 and demonstrate fine structure of the soft diffuse emission, contrasting with relatively smooth and well-confined hard emission.
For better characterization of the diffuse emission, we identified bright point sources and masked them along with the central source SS 433. The resulting composite map of the residual diffuse emission is shown in Figure 5. The overlaid distance and angle rulers highlight the characteristic dimensions and morphology of the prominent structures. Namely, one can see that the “Extended X-ray Jets” start pc away from the central source and demonstrate remarkable symmetry, with the Eastern one starting only a couple of parsecs further from the central source. Both EXJs, characterized by hard X-ray spectra, terminate around pc. This symmetry contrasts with the more asymmetric and irregular soft emission of the lobes.
The same conclusion is even better illustrated by the side-by-side comparison of the two lobes in Fig. 4. Indeed, while the Eastern lobe extends up to , the Western one terminates at pc. The inclination of the relativistic jet’s precession axis to the line of sight is (the relativistic jet on the eastern side is on average approaching us), so this difference cannot be explained by geometrical effects only. However, it broadly agrees with the assumption that the energy flow from the central source interacts with the denser ambient medium in the western part, reflecting the gas density gradient perpendicular to the equatorial plane of the Galactic disk (e.g., Goodall et al., 2011a). Since the difference between the lobes’ dimensions amounts to a sizeable fraction of them, one can infer the exponential scale-height of the ambient gas density profile pc, consistent with the scale-height of the Galactic neutral gas disk. The X-ray emission from the Northern and Southern boundaries of the nebula appears soft and faint, being noticeably affected by the foreground dust absorption, especially in their Western parts (as illustrated in Appendix C).
Taking this into account, we conclude that the soft X-ray emission can be traced across the full extent of the nebula, except for the “inner cavity” with a radius of pc. The outer boundary of the emission of the Northern and Southern arcs (outlined as dashed annuli including Region 4 in Fig. 5) is located at from the center, coinciding with the boundary of the W50’s radio emission. This morphology suggests that the soft X-ray emitting gas is confined in a thick () close-to-spherical shell, driving a strong shock in the surrounding ISM.
The Eastern and the Western lobes are roughly confined inside the projection of the cone with the half-opening angle , while the EXJs encompass only the cone with a twice smaller opening angle,. From the eROSITA spectrally decomposed RGB images, a gentle spectral evolution of the EXJs emission with distance is seen clearly. EXJ emission is significantly harder at its sharp inner boundary than further away from SS 433, where EXJs blend with softer thermal emission that fills the rest of the lobe, in agreement with earlier findings with XMM-Newton and NuSTAR (Brinkmann et al., 2007; Safi-Harb et al., 2022).
5 X-ray spectra
5.1 Central source
The spectral capabilities of eROSITA have been demonstrated using many objects, including SS 433 itself (e.g., Sunyaev et al., 2021). Since this source is extremely bright, the observations were intentionally performed during the two consecutive orbital eclipses of the compact object (and the jets) by the companion star, so its soft X-ray flux was dimmer by a factor of several compared to the typical levels (Kawai et al., 1989; Lopez et al., 2006; Marshall et al., 2013). The combined X-ray spectrum of SS 433 extracted from arcmin aperture (with background estimated from to 5 arcmin annuli, being however completely negligible compared the source) using the data of all pointings in which it falls within Field-of-View is shown in Fig. 6.
As expected, it is well described by the model of the optically thin multitemperature emission from collimated flows that adiabatically expand and cool at distances from a few to a few cm from the central object (Brinkmann et al., 1988, 1991, 2005; Kotani et al., 1996; Marshall et al., 2002; Medvedev & Fabrika, 2010; Khabibullin et al., 2016; Medvedev et al., 2019). We discuss it more in Appendix A, while the main focus of this study is instead on the much fainter diffuse emission from the surrounding nebula.
5.2 Diffuse emission
The high sensitivity of the presented data allowed us to perform spectroscopy of the diffuse emission in several regions and determine the physical state of the radiating plasma. We have selected eight regions, shown as numbered boxes in Figure 5, for detailed spectral analysis of their X-ray emission. For each region, the background (estimated from the adjacent regions) was subtracted to isolate the net signal. We present the parameters of the fitted models in Table 1 and discuss the implications of the corresponding results below.
5.2.1 Eastern Lobe
Emission from the brightest filament in the Eastern lobe (Region 1 in Figure 5) features a spectrum characteristic for a thermal (but non-equilibrium) optically-thin plasma with a temperature of , as evidenced by the line emission from the highly ionized oxygen, neon, magnesium, silicon and iron (see the left panel in Figure 7). The line emissivity and ratios are consistent with the Solar abundance of heavy elements, but also indicate a non-equilibrium state of the gas ionization. Namely, it is qualitatively consistent with the expectation for a plane-parallel shell of the gas that experienced a passage of the shock with velocity (from the Rankine-Hugoniot relations for the temperature and density jump properties) some yr ago.
The inferred hydrogen number density, cm-3, is a factor of a few higher than the expected gas density of the interstellar gas pc away from the plane of the Galactic disk. The spectrum of the diffuse emission in the Eastern lobe (exemplified by Region 3 in Figure 5) also can be well described by thermal non-equilibrium emission of the gas with somewhat higher temperature keV and lower hydrogen number density cm-3 (see the left panel in Figure 7).
5.2.2 Western Lobe
The emission from the Western lobe appears to be noticeably harder, partially due to the higher interstellar absorption in that direction. However, it also corresponds to the emission of plasma with genuinely higher temperature, keV, and a similar hydrogen number density, cm-3 (and a few times higher for the brightest filament), as illustrated by spectra from Regions 6 and 7 in Figure 5). The non-equilibrium ionization model (see the right panel in Figure 7) results in a comparable estimate of yr for the ionization time-scale.
The total absorption-corrected luminosity of the diffuse soft X-ray emission from the nebula is in 0.5-4 keV, which is by a factor of few tens smaller than the apparent X-ray luminosity of the central object, few ,222Due to the beaming of the radiation along the thick accretion disc axis (expected in the case of highly supercritical accretion), the actual X-ray luminosity of SS 433 might be orders of magnitude larger, namely comparable to the Eddington luminosity and kinetic luminosity of the jets, (e.g. Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007; Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016; Middleton et al., 2021; Fogantini et al., 2023). and much smaller than the supposed kinetic luminosity of the outflows from the supercritical accretion disk, . However, the thermal energy content of the X-ray emitting gas in the Eastern lobe (with estimated gas number density 0.1 cm-3, temperature 0.5 keV and volume assuming roughly to cylindrical geometry ) is erg, requiring a supply of at least over years. The mass of this gas amounts to , implying that the bulk of it cannot be provided by the outflows from SS 433. If the volume filling factor of the emitting medium is, however, low, the estimate for the required mass and energy could be decreased by a factor of , but similarly, the density should be increased, leading to the corresponding decrease in the time since the shock passage. That means that the X-ray emitting gas provides an estimate of the average energy input in the lobes at the level of , which is comparable to the kinetic luminosity of the jets and the Eddington luminosity of the system (for the mass of the compact object below 10).
A similar estimate can be obtained from the Western lobe, implying nearly symmetric energy injection along the nebula’s axis at a very high rate over the last few tens of thousands of years. Given that the temperature inferred from the spectral fits with the non-equilbirum emission is driven mostly by characteristics of the ionization state, it is likely a lower limit on the actual temperature of the emitting gas. Hence, a significant portion of the gas energy content might be effectively hidden, while the required energy input should be increased correspondingly.
5.2.3 Northern and southern boundaries
The emission from the nebula boundaries is exemplified by the Northern Arc region and the Southern Extension (Regions 4 and 8 in Figure 5). The thermal non-equilibrium model provides the temperature estimate keV, low hydrogen number density, cm-3 for the Northern Arc and even lower for the Southern Extension. The ionization timescale is, however, similar here as well, yr. This observation contrasts with the brightness of the radio emission, which in the Northern Arc region is one of the brightest across the nebula. The high degree of linear polarization of the radio emission might indicate that the preferential orientation of the interstellar magnetic field in this region favors efficient acceleration of relativistic particles producing synchrotron emission. We note here that the central part of the W50 X-ray map suffers from higher photoelectric absorption than the Eastern and even Western lobes, as illustrated in Fig. 9. This excess absorption is likely associated with the foreground gas and has no physical connection to SS 433/W50.
5.2.4 Extended X-ray jets
In contrast to this “thermal” picture, the X-ray spectra of X-ray emission from the “extended X-ray jets” demonstrate a featureless spectrum that can be described by a power law with the slope in the range from (at their base, e.g., Region 5 in Fig. 5) to (at the termination points, e.g. Region 2 in Fig. 5). In agreement with previous spectral studies of this emission (e.g., Brinkmann et al., 2007), the thermal bremsstrahlung model also provides an equally good fit with temperatures from keV (at the termination) to keV (at the base). The gas emissivity required to produce the X-ray luminosity at level implies hydrogen number densities cm-3 in these regions (assuming a filling factor of unity). Clearly, if the thermal nature of this emission were true, it would have a drastically higher pressure compared to the ambient soft X-ray emitting gas and would expand sideways supersonically, driving a strong shock wave. Alternatively, the emission might result from the population of relativistic particles accelerated within EXJs. The detection of polarization in X-rays (Kaaret et al., 2024) and TeV emission from regions cospatial with the EXJs (Abeysekara et al., 2018; Alfaro et al., 2024; H. E. S. S. Collaboration et al., 2024) strongly supports the non-thermal scenario.
The absence of any noticeable radio, infrared, or optical emission from these structures would then require substantial self-absorption of radiation at these wavelengths (which is unlikely given that the gas number density is likely very low inside them) or a rather peculiar energy distribution of the relativistic particles with a large fraction of them having energies in excess of tens of TeV. Another possibility could be scattering of very bright and beamed emission of the central source (e.g., Panferov & Kabanov, 2013; Khabibullin & Sazonov, 2016, 2019), but the small optical depth with respect to the Thomson scattering, , and the observed spectral gradient along these structures disfavor this possibility.
6 Discussion
The Eastern EXJ was well mapped in X-rays with XMM-Newton, Chandra, and NuSTAR (e.g., Brinkmann et al., 2007; Safi-Harb et al., 2022), revealing its non-thermal nature. Furthermore, recent IXPE observation (Kaaret et al., 2024) revealed polarized X-ray emission from a patch adjacent to the base of the Eastern EXJ. These data suggest that X-ray photons are produced by a synchrotron mechanism with the magnetic field predominantly aligned with the EXJ axis. The eROSITA data show that the bases of both EXJs are located at essentially the same distances from SS 433, have similarly sharp inner boundaries, and non-thermal (featureless) spectra across their entire extents.
These structures have often been considered as a direct consequence of the relativistic jets’ propagation inside the nebula before they deposit the bulk of their kinetic energy in the lobes, energizing them and the whole extent of the nebula (e.g., Panferov, 2014). However, the direction of the SS 433 jets is known to precess about the mean axis with the amplitude degrees, which is a factor of 2 larger than the observed opening angle of EXJ. Since this extended X-ray emission comes from distances of tens of pc, i.e. orders of magnitude further away from the central source than the compact radio and X-ray jets, which are traced up to a few cm, the actual connection between them is not obvious (e.g., Panferov, 2017). It might be governed by recollimation of the jets’ flows at some intermediate distance from the source (Eichler, 1983; Peter & Eichler, 1993), a historical change in the jets’ direction (Goodall et al., 2011a; Monceau-Baroux et al., 2015), or a different regime of baryonic jets propagation through the magnetized ISM after their gas “freezes out” after recombining due to rapid adiabatic expansion, cooling losses and lack of ambient interactions (Churazov et al., 2020).
If we assume that there were no recent drastic changes in the properties of W50/SS 433 (e.g., temporary quenching of the central source), the overall picture of the “large-scale energy flow” in the W50/SS 433, supported by X-ray data, can be schematically decomposed into three major components (illustrated in Fig. 8):
“Dark flow”: from a compact source up to the base of EXJs covering the range of distances from to from SS 433. Over this distance range, there are no visible signatures in the X-ray (or any other) band.
“Non-thermal flow”: from the base to the end of EXJs that have featureless non-thermal spectra (presumably, synchrotron emission).
“Thermal flow”: extended diffuse regions between the W50 radio boundaries and EXJs. Its X-ray emission is typical for ISM shock-heated to sub-keV temperatures.
Analytical and numerical models primarily aimed at reproducing W50’s radio morphology often assume that a quasi-spherical part of the nebula is a remnant of the supernova explosion, while the elongated structures are directly associated with the energy release by the binary system (e.g., Begelman et al., 1980; Zealey et al., 1980; Eichler, 1983; Peter & Eichler, 1993; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Asahina et al., 2014; Monceau-Baroux et al., 2015; Panferov, 2017; Ohmura et al., 2021). In the model considered by Churazov et al. (2024), the entire nebula is powered by the central source. There, an anisotropic outflow/wind from the Hyper-Eddington accretor is assumed - a combination of a more powerful collimated outflow along the orbital axis and a quasi-isotropic wind in all other directions. In this model, the “Dark flow” phase corresponds to the free expansion of the wind, and the onset of the EXJs is associated with the termination shock of the isotropic wind, which initiates shocks in the collimated flow where the pressure rises suddenly. Finally, the thermal part is the shock-heated ISM. In this model, the characteristic shape of the W50 nebula (a quasi-spherical part and two extensions) reflects the anisotropy of the velocity and kinetic power of the wind, with no need for the impact of the narrow baryonic jets at all.
In all these models, the outer shock, which delineates the radio boundary of W50, is not dissimilar to the shocks seen around middle-aged supernova remnants or stellar wind-blown bubbles (e.g., Konigl, 1983; Dubner et al., 1998). What makes W50 “special” is the presence of EXJs, which are radio-faint but bright in X-rays and VHE gamma-rays. These structures are also plausibly related to efficient particle acceleration at strong shocks, but with a shock configuration that is rather different compared to typical SNR shocks (e.g. Bykov et al., 2025). Remarkable symmetry of the EXJs, very clearly demonstrated by the eROSITA image, points towards irrelevance of the ambient ISM properties to the energy and momentum propagation, as well as the particle acceleration regime, meaning that a cocoon or a wind-blown cavity likely encompasses these structures in the current epoch. In this regard, the presence of H emitting filaments in several places across the nebula might help probe the energetic content of the nebula (e.g., Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017). Interestingly, the location of the brightest H filaments does not correspond to any prominent features in thermal X-ray emission. This might be the case when the H emission comes from the outer boundary of the nebula, where the shock wave is already in the radiative mode. Significantly higher gas temperature obtained for the emission from the Western lobe might also be indicative of this process, namely that the radiative shock wave runs against the Galactic density gradient and stalls earlier in this direction, evacuating a smaller volume to be filled by the hot X-ray gas. Given the presence of bright silicon lines in the spectra extracted from these regions, an observation of them with the XRISM observatory (Tashiro et al., 2025) might provide estimates of the directed and turbulent velocities. A large grasp soft X-ray microcalorimeter mission, like LEM (Kraft et al., 2022) or HUBS (Bregman et al., 2023) would also be perfectly suited for in-depth spectral exploration of this complex system with multiple interacting components.
7 Conclusions
SRG/eROSITA observations of W50/SS 433 provide, for the first time, the entire X-ray map of the nebula with high spatial and spectral resolution. The data support a physical picture in which the anisotropic energy flow from SS 433 evolves passing through three distinct stages:
an invisible “dark” flow between 0.1 and 25 pc (=anisotropic wind from the binary system)
a “non-thermal” flow over another (=EXJs)
a thermal flow (= shock-heated ISM) that envelopes EXJs.
The appearance of the nebula is further affected by the large ambient density gradients and a heavy foreground photoelectric absorption that projects almost exactly on the central quasi-spherical part of the W50 nebula.
The thermal part of the W50 X-ray emission can be reasonably well described by the shock-heated plasma that has not yet reached temperature and ionization equilibrium. Such emission is typical for middle-age or old SNRs. The outer radio boundary of the nebula is also reminiscent of the SNR shocks.
On the contrary, the “Extended X-ray Jets” are the most remarkable features of this system on scales of tens of pc. Their sharp inner edges plausibly correspond to extreme shocks that accelerate particles and power the X-ray (synchrotron) emission, and, at 10 orders of magnitude higher energies, the TeV emission. W50/SS 433 system illustrates clearly the important role the Hyper-Eddington accretors might play for the energetics of the ISM in galaxies at different redshifts and the production of ultra-high-energy particles.
Acknowledgements.
This work is partly based on observations with the eROSITA telescope onboard SRG space observatory. The SRG observatory was built by Roskosmos in the interests of the Russian Academy of Sciences, represented by its Space Research Institute (IKI) in the framework of the Russian Federal Space Program, with the participation of the Deutsches Zentrum für Luft- und Raumfahrt (DLR). The eROSITA X-ray telescope was built by a consortium of German Institutes led by MPE, and supported by DLR. The SRG spacecraft was designed, built, launched, and operated by the Lavochkin Association and its subcontractors. The science data are downlinked via the Deep Space Network Antennae in Bear Lakes, Ussurijsk, and Baikonur, funded by Roskosmos. The development and construction of the eROSITA X-ray instrument was led by MPE, with contributions from the Dr. Karl Remeis Observatory Bamberg ECAP (FAU Erlangen-Nuernberg), the University of Hamburg Observatory, the Leibniz Institute for Astrophysics Potsdam (AIP), and the Institute for Astronomy and Astrophysics of the University of Tübingen, with the support of DLR and the Max Planck Society. The Argelander Institute for Astronomy of the University of Bonn and the Ludwig Maximilians Universität München also participated in the science preparation for eROSITA. The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium. IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.References
- Abell & Margon (1979) Abell, G. O. & Margon, B. 1979, Nature, 279, 701
- Abeysekara et al. (2018) Abeysekara, A. U., Albert, A., Alfaro, R., et al. 2018, Nature, 562, 82
- Abolmasov et al. (2010) Abolmasov, P., Maryeva, O., & Burenkov, A. N. 2010, AN, 331, 412
- Alfaro et al. (2024) Alfaro, R., Alvarez, C., Arteaga-Velázquez, J. C., et al. 2024, ApJ, 976, 30
- Arnason et al. (2021) Arnason, R. M., Papei, H., Barmby, P., Bahramian, A., & Gorski, M. D. 2021, MNRAS, 502, 5455
- Arnaud (1996) Arnaud, K. A. 1996, in Astronomical Society of the Pacific Conference Series, Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes, 17
- Asahina et al. (2014) Asahina, Y., Ogawa, T., Kawashima, T., et al. 2014, ApJ, 789, 79
- Atapin & Fabrika (2016) Atapin, K. E. & Fabrika, S. N. 2016, AstL, 42, 517
- Begelman et al. (2006) Begelman, M. C., King, A. R., & Pringle, J. E. 2006, MNRAS, 370, 399
- Begelman et al. (1980) Begelman, M. C., Sarazin, C. L., Hatchett, S. P., McKee, C. F., & Arons, J. 1980, ApJ, 238, 722
- Blundell & Bowler (2004) Blundell, K. M. & Bowler, M. G. 2004, ApJL, 616, L159
- Blundell et al. (2011) Blundell, K. M., Schmidtobreick, L., & Trushkin, S. 2011, MNRAS, 417, 2401
- Borkowski et al. (2001) Borkowski, K. J., Lyerly, W. J., & Reynolds, S. P. 2001, ApJ, 548, 820
- Boumis et al. (2007) Boumis, P., Meaburn, J., Alikakos, J., et al. 2007, MNRAS, 381, 308
- Bregman et al. (2023) Bregman, J., Cen, R., Chen, Y., et al. 2023, SCPMA, 66, 299513
- Brinkmann et al. (1996) Brinkmann, W., Aschenbach, B., & Kawai, N. 1996, A&A, 312, 306
- Brinkmann et al. (1988) Brinkmann, W., Fink, H. H., Massaglia, S., Bodo, G., & Ferrari, A. 1988, A&A, 196, 313
- Brinkmann & Kawai (2000) Brinkmann, W. & Kawai, N. 2000, A&A, 363, 640
- Brinkmann et al. (1991) Brinkmann, W., Kawai, N., Matsuoka, M., & Fink, H. H. 1991, A&A, 241, 112
- Brinkmann et al. (2005) Brinkmann, W., Kotani, T., & Kawai, N. 2005, A&A, 431, 575
- Brinkmann et al. (2007) Brinkmann, W., Pratt, G. W., Rohr, S., Kawai, N., & Burwitz, V. 2007, A&A, 463, 611
- Broderick et al. (2018) Broderick, J. W., Fender, R. P., Miller-Jones, J. C. A., et al. 2018, MNRAS, 475, 5360
- Bykov et al. (2025) Bykov, A. M., Osipov, S. M., Romansky, V. I., et al. 2025, PhRvD, 112, 063017
- Cherepashchuk et al. (2023) Cherepashchuk, A., Belinski, A., Dodin, A., & Postnov, K. 2023, NewA, 103, 102060
- Cherepashchuk (1981) Cherepashchuk, A. M. 1981, MNRAS, 194, 761
- Cherepashchuk et al. (2025) Cherepashchuk, A. M., Dodin, A. V., & Postnov, K. A. 2025, arXiv, arXiv:2506.01106
- Cherepashchuk et al. (2018) Cherepashchuk, A. M., Esipov, V. F., Dodin, A. V., Davydov, V. V., & Belinskii, A. A. 2018, ARep, 62, 747
- Chi et al. (2024) Chi, Y.-H., Huang, J., Zhou, P., et al. 2024, ApJL, 975, L28
- Churazov et al. (2023) Churazov, E., Khabibullin, I., Bykov, A. M., Lyskova, N., & Sunyaev, R. 2023, A&A, 670, A156
- Churazov et al. (2021a) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021a, A&A, 651, A41
- Churazov et al. (2025) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Dolag, K. 2025, arXiv, arXiv:2507.19987
- Churazov et al. (2020) Churazov, E., Khabibullin, I., & Sunyaev, R. 2020, MNRAS, 495, L51
- Churazov et al. (2024) Churazov, E. M., Khabibullin, I. I., & Bykov, A. M. 2024, A&A, 688, A4
- Churazov et al. (2021b) Churazov, E. M., Khabibullin, I. I., Bykov, A. M., et al. 2021b, MNRAS, 507, 971
- Dubner et al. (1998) Dubner, G. M., Holdaway, M., Goss, W. M., & Mirabel, I. F. 1998, AJ, 116, 1842
- Eichler (1983) Eichler, D. 1983, ApJ, 272, 48
- Eikenberry et al. (2001) Eikenberry, S. S., Cameron, P. B., Fierce, B. W., et al. 2001, ApJ, 561, 1027
- Elston & Baum (1987) Elston, R. & Baum, S. 1987, AJ, 94, 1633
- Fabian & Rees (1979) Fabian, A. C. & Rees, M. J. 1979, MNRAS, 187, 13P
- Fabrika (2004) Fabrika, S. 2004, ASPRv, 12, 1
- Fabrika & Mescheryakov (2001) Fabrika, S. & Mescheryakov, A. 2001, in IAU Symposium, Vol. 205, Galaxies and their Constituents at the Highest Angular Resolutions, ed. R. T. Schilizzi, 268
- Farnes et al. (2017) Farnes, J. S., Gaensler, B. M., Purcell, C., et al. 2017, MNRAS, 467, 4777
- Filippova et al. (2006) Filippova, E., Revnivtsev, M., Fabrika, S., Postnov, K., & Seifina, E. 2006, A&A, 460, 125
- Fogantini et al. (2023) Fogantini, F. A., García, F., Combi, J. A., et al. 2023, A&A, 669, A149
- Forman et al. (1978) Forman, W., Jones, C., Cominsky, L., et al. 1978, ApJS, 38, 357
- Fragile et al. (2025) Fragile, P. C., Middleton, M. J., Bollimpalli, D. A., & Smith, Z. 2025, MNRAS, 540, 2820
- Fuchs et al. (2006) Fuchs, Y., Koch Miramond, L., & Ábrahám, P. 2006, A&A, 445, 1041
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Geldzahler et al. (1980) Geldzahler, B. J., Pauls, T., & Salter, C. J. 1980, A&A, 84, 237
- Goodall et al. (2011a) Goodall, P. T., Alouani-Bibi, F., & Blundell, K. M. 2011a, MNRAS, 414, 2838
- Goodall et al. (2011b) Goodall, P. T., Blundell, K. M., & Bell Burnell, S. J. 2011b, MNRAS, 414, 2828
- Goranskij (2011) Goranskij, V. 2011, PZ, 31, 5
- Green (2025) Green, D. A. 2025, JApA, 46, 14
- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
- Gregory et al. (1996) Gregory, P. C., Scott, W. K., Douglas, K., & Condon, J. J. 1996, ApJS, 103, 427
- Grindlay et al. (1984) Grindlay, J. E., Band, D., Seward, F., et al. 1984, ApJ, 277, 286
- H. E. S. S. Collaboration et al. (2024) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2024, Sci, 383, 402
- Heinz & Sunyaev (2002) Heinz, S. & Sunyaev, R. 2002, A&A, 390, 751
- Hjellming & Johnston (1981a) Hjellming, R. M. & Johnston, K. J. 1981a, ApJL, 246, L141
- Hjellming & Johnston (1981b) Hjellming, R. M. & Johnston, K. J. 1981b, Nature, 290, 100
- Holden & Caswell (1969) Holden, D. J. & Caswell, J. L. 1969, MNRAS, 143, 407
- Jiang et al. (2019) Jiang, Y.-F., Stone, J. M., & Davis, S. W. 2019, ApJ, 880, 67
- Kaaret et al. (2017) Kaaret, P., Feng, H., & Roberts, T. P. 2017, ARA&A, 55, 303
- Kaaret et al. (2024) Kaaret, P., Ferrazzoli, R., Silvestri, S., et al. 2024, ApJL, 961, L12
- Katz (1980) Katz, J. I. 1980, ApJL, 236, L127
- Katz et al. (1982) Katz, J. I., Anderson, S. F., Margon, B., & Grandi, S. A. 1982, ApJ, 260, 780
- Kawai et al. (1989) Kawai, N., Matsuoka, M., Pan, H.-C., & Stewart, G. C. 1989, PASJ, 41, 491
- Khabibullin et al. (2022) Khabibullin, I., Churazov, E., & Sunyaev, R. 2022, MNRAS, 509, 6068
- Khabibullin et al. (2016) Khabibullin, I., Medvedev, P., & Sazonov, S. 2016, MNRAS, 455, 1414
- Khabibullin & Sazonov (2016) Khabibullin, I. & Sazonov, S. 2016, MNRAS, 457, 3963
- Khabibullin et al. (2023) Khabibullin, I. I., Churazov, E. M., Bykov, A. M., Chugai, N. N., & Sunyaev, R. A. 2023, MNRAS, 521, 5536
- Khabibullin et al. (2024) Khabibullin, I. I., Churazov, E. M., Chugai, N. N., et al. 2024, A&A, 689, A278
- Khabibullin & Sazonov (2012) Khabibullin, I. I. & Sazonov, S. Y. 2012, AstL, 38, 443
- Khabibullin & Sazonov (2017) Khabibullin, I. I. & Sazonov, S. Y. 2017, AstL, 43, 388
- Khabibullin & Sazonov (2019) Khabibullin, I. I. & Sazonov, S. Y. 2019, AstL, 45, 282
- Kimura et al. (2020) Kimura, S. S., Murase, K., & Mészáros, P. 2020, ApJ, 904, 188
- Kirshner & Chevalier (1980) Kirshner, R. P. & Chevalier, R. A. 1980, ApJL, 242, L77
- Konigl (1983) Konigl, A. 1983, MNRAS, 205, 471
- Kotani et al. (1996) Kotani, T., Kawai, N., Matsuoka, M., & Brinkmann, W. 1996, PASJ, 48, 619
- Kraft et al. (2022) Kraft, R., Markevitch, M., Kilbourne, C., et al. 2022, arXiv, arXiv:2211.09827
- Kubota et al. (2010) Kubota, K., Ueda, Y., Kawai, N., et al. 2010, PASJ, 62, 323
- Leibowitz (1984) Leibowitz, E. M. 1984, MNRAS, 210, 279
- LHAASO Collaboration (2024) LHAASO Collaboration. 2024, arXiv, arXiv:2410.08988
- Lopez et al. (2006) Lopez, L. A., Marshall, H. L., Canizares, C. R., Schulz, N. S., & Kane, J. F. 2006, ApJ, 650, 338
- Margon (1984) Margon, B. 1984, ARA&A, 22, 507
- Margon et al. (1979) Margon, B., Ford, H. C., Katz, J. I., et al. 1979, ApJL, 230, L41
- Marshall et al. (2013) Marshall, H. L., Canizares, C. R., Hillwig, T., et al. 2013, ApJ, 775, 75
- Marshall et al. (2002) Marshall, H. L., Canizares, C. R., & Schulz, N. S. 2002, ApJ, 564, 941
- Medvedev & Fabrika (2010) Medvedev, A. & Fabrika, S. 2010, MNRAS, 402, 479
- Medvedev et al. (2019) Medvedev, P. S., Khabibullin, I. I., & Sazonov, S. Y. 2019, AstL, 45, 299
- Medvedev et al. (2018) Medvedev, P. S., Khabibullin, I. I., Sazonov, S. Y., Churazov, E. M., & Tsygankov, S. S. 2018, AstL, 44, 390
- Medvedev et al. (2022) Medvedev, P. S., Khabibullin, I. I., Semena, A. N., et al. 2022, AstL, 48, 389
- Middleton et al. (2021) Middleton, M. J., Walton, D. J., Alston, W., et al. 2021, MNRAS, 506, 1045
- Migliari et al. (2005) Migliari, S., Fender, R. P., Blundell, K. M., Méndez, M., & van der Klis, M. 2005, MNRAS, 358, 860
- Miller-Jones et al. (2008) Miller-Jones, J. C. A., Migliari, S., Fender, R. P., et al. 2008, ApJ, 682, 1141
- Monceau-Baroux et al. (2015) Monceau-Baroux, R., Porth, O., Meliani, Z., & Keppens, R. 2015, A&A, 574, A143
- Namiki et al. (2000) Namiki, M., Kawai, N., Kotani, T., Mamauchi, S., & Brinkmann, W. 2000, AdSpR, 25, 709
- Ohmura et al. (2021) Ohmura, T., Ono, K., Sakemi, H., et al. 2021, ApJ, 910, 149
- Ohsuga & Mineshige (2011) Ohsuga, K. & Mineshige, S. 2011, ApJ, 736, 2
- Panferov (2014) Panferov, A. 2014, A&A, 562, A130
- Panferov (2017) Panferov, A. A. 2017, A&A, 599, A77
- Panferov & Kabanov (2013) Panferov, A. A. & Kabanov, A. A. 2013, arXiv, arXiv:1306.4486
- Peretti et al. (2025) Peretti, E., Petropoulou, M., Vasilopoulos, G., & Gabici, S. 2025, A&A, 698, A188
- Peter & Eichler (1993) Peter, W. & Eichler, D. 1993, ApJ, 417, 170
- Poutanen et al. (2007) Poutanen, J., Lipunova, G., Fabrika, S., Butkevich, A. G., & Abolmasov, P. 2007, MNRAS, 377, 1187
- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
- Predehl et al. (2020) Predehl, P., Sunyaev, R. A., Becker, W., et al. 2020, Nature, 588, 227
- Rosado et al. (2021) Rosado, M., Sánchez-Cruces, M., Ambrocio-Cruz, P., & Trejo, D. 2021, MNRAS, 506, 4263
- Safi-Harb et al. (2022) Safi-Harb, S., Mac Intyre, B., Zhang, S., et al. 2022, ApJ, 935, 163
- Safi-Harb & Ögelman (1997) Safi-Harb, S. & Ögelman, H. 1997, ApJ, 483, 868
- Safi-Harb & Petre (1999) Safi-Harb, S. & Petre, R. 1999, ApJ, 512, 784
- Sakai et al. (2025) Sakai, Y., Yamada, S., Sakemi, H., et al. 2025, arXiv, arXiv:2507.19042
- Sakemi et al. (2021) Sakemi, H., Omae, R., Ohmura, T., & Machida, M. 2021, PASJ, 73, 530
- Sarazin et al. (1980) Sarazin, C. L., Begelman, M. C., & Hatchett, S. P. 1980, ApJL, 238, L129
- Seward et al. (1980) Seward, F., Grindlay, J., Seaquist, E., & Gilmore, W. 1980, Nature, 287, 806
- Seward et al. (1976) Seward, F. D., Page, C. G., Turner, M. J. L., & Pounds, K. A. 1976, MNRAS, 175, 39P
- Shakura & Sunyaev (1973) Shakura, N. I. & Sunyaev, R. A. 1973, A&A, 24, 337
- Sharpless (1959) Sharpless, S. 1959, ApJS, 4, 257
- Shklovskii (1981) Shklovskii, I. S. 1981, SvA, 25, 315
- Shuder et al. (1980) Shuder, J. M., Hatfield, B. F., & Cohen, R. D. 1980, PASP, 92, 259
- Spencer (1979) Spencer, R. E. 1979, Nature, 282, 483
- Stephenson & Sanduleak (1977) Stephenson, C. B. & Sanduleak, N. 1977, ApJS, 33, 459
- Stewart et al. (1987) Stewart, G. C., Watson, M. G., Matsuoka, M., et al. 1987, MNRAS, 228, 293
- Sunyaev et al. (2021) Sunyaev, R., Arefiev, V., Babyshkin, V., et al. 2021, A&A, 656, A132
- Tashiro et al. (2025) Tashiro, M., Kelley, R., Watanabe, S., et al. 2025, PASJ
- Tsuji et al. (2025) Tsuji, N., Inoue, Y., Khangulyan, D., et al. 2025, arXiv, arXiv:2510.06431
- van den Heuvel (1981) van den Heuvel, E. P. J. 1981, VA, 25, 95
- van den Heuvel et al. (1980) van den Heuvel, E. P. J., Ostriker, J. P., & Petterson, J. A. 1980, A&A, 81, L7
- van den Heuvel et al. (2017) van den Heuvel, E. P. J., Portegies Zwart, S. F., & de Mink, S. E. 2017, MNRAS, 471, 4256
- Velázquez & Raga (2000) Velázquez, P. F. & Raga, A. C. 2000, A&A, 362, 780
- Wang et al. (2025) Wang, J., Reville, B., & Aharonian, F. A. 2025, ApJL, 989, L25
- Watson et al. (1986) Watson, M. G., Stewart, G. C., Brinkmann, W., & King, A. R. 1986, MNRAS, 222, 261
- Watson et al. (1983) Watson, M. G., Willingale, R., Grindlay, J. E., & Seward, F. D. 1983, ApJ, 273, 688
- Westerhout (1958) Westerhout, G. 1958, BAN, 14, 215
- Whitmire & Matese (1980) Whitmire, D. P. & Matese, J. J. 1980, MNRAS, 193, 707
- Wilms et al. (2000) Wilms, J., Allen, A., & McCray, R. 2000, ApJ, 542, 914
- Yamauchi et al. (1994) Yamauchi, S., Kawai, N., & Aoki, T. 1994, PASJ, 46, L109
- Yoshioka et al. (2022) Yoshioka, S., Mineshige, S., Ohsuga, K., Kawashima, T., & Kitaki, T. 2022, PASJ, 74, 1378
- Zavala et al. (2008) Zavala, J., Velázquez, P. F., Cerqueira, A. H., & Dubner, G. M. 2008, MNRAS, 387, 839
- Zealey et al. (1980) Zealey, W. J., Dopita, M. A., & Malin, D. F. 1980, MNRAS, 192, 731
Appendix A Spectrum of the central source
As one can see in Fig. 6, both blue-shifted and red-shifted lines of highly ionized (H-like and He-like) heavy elements (in particular, silicon, sulfur, iron, and nickel) moving at 0.26 speed of light can be easily discriminated and identified. They arise from the approaching (with apparent redshift -0.085) and the receding jet (with apparent redshift 0.167), respectively. In agreement with the findings of numerous previous observations (see Medvedev et al. 2019, for a recent summary), a factor of 10 overabundance of nickel (with respect to iron) in the gas of the jets is required to describe the spectrum well (most clearly indicated by the ratio of the Ni XXVII and Fe XXV lines). Spectral resolution of eROSITA is not sufficient to resolve the width of the lines, which is known to be dominated by the ballistic motion of gas within a cone of with opening angle , as measured by high resolution spectroscopy with Chandra gratings (e.g. Marshall et al. 2002). Some residual emission can also be observed in the iron K complex region between 6 and 7 keV, which could come from the reflection of central source emission on the walls of the optically thick accretion disk funnel (Medvedev & Fabrika 2010) or in the optically thin accretion disk wind (Medvedev et al. 2018, 2019). Comparison of the spectra taken in the two epochs of observations separated by one orbital period shows very good consistency between them333Since one orbital period separation is small compared to the precession period and the observations were close to an extreme point of the precession curve, the change in the Doppler shifts of the jets is small., with the biggest difference observed in the 6-7 keV band. This might be related to the change in the visible geometry of this region or to the physical conditions within it (see a discussion in Medvedev et al. 2019). High-resolution spectroscopy of this emission with the XRISM observatory (Tashiro et al. 2025) will open new possibilities for characterizing this emission and potentially give access to more intricate diagnostics based on velocity structure and resonant scattering in the jets (Khabibullin & Sazonov 2012), as well as possible signatures of multiphase cooling and ionization state dynamics in supersonically expanding jets prone to thermal instability and fragmentation (Brinkmann et al. 1988; Brinkmann & Kawai 2000; Khabibullin et al. 2016). The SS433’s spectrum presented here illustrates the spectroscopic capabilities of eROSITA and shows that, for sufficiently bright sources, reliable continuum and line characterization can be achieved up to 10 keV.
Appendix B Parameters of the spectral models
Here we list parameters of the best-fit models used to describe spectra shown in Fig. 7 and discussed in Sec. 5. The spectral analysis was performed using standard routines of the XSPEC package (Arnaud 1996). Namely, the absorbed (using tbabs model by Wilms et al. 2000) non-equilibrium plasma emission model (nei model by Borkowski et al. 2001) was used to describe thermal emission, while similarly absorbed thermal bremsstrahlung model (tbabs*brems in XSPEC notation) was used for a phenomenological description of the non-thermal emission from the EXJs. Due to the complexity of the system, the X-ray emission from all these regions is likely a mixture of several components; hence, one needs to take these derived parameters with caution and probably as approximate guidelines for the physical conditions within them. For the background estimation, adjacent regions of significant size were used, and the signal inside them was subtracted from the corresponding source region. For all regions except for the relatively faint Regions 4 and 8, the relative contribution of the background emission in the source regions was rather small.
| # | Norm | |||
| keV | ||||
| Thermal regions (tbabs*vnei) | ||||
| 1 | 0.670.02 | 0.290.02 | 0.260.05 | 12.333.63 |
| 3 | 0.560.02 | 0.560.07 | 0.110.01 | 0.800.19 |
| 4 | 0.790.05 | 0.950.25 | 0.090.01 | 0.210.07 |
| 6 | 0.850.03 | 0.760.06 | 0.640.12 | 2.010.28 |
| 7 | 0.720.02 | 1.300.18 | 0.080.01 | 0.880.13 |
| 8 | 0.750.05 | 0.780.22 | 0.130.05 | 0.420.18 |
| Non-thermal regions (tbabs*brems) | ||||
| 2 | 0.510.01 | 3.050.15 | - | 1.620.04 |
| 5 | 0.840.03 | 14.53.9 | - | 0.640.01 |
Appendix C Impact of foreground absorption
In this section, we demonstrate that some of the large-scale variations in the diffuse X-ray emission from W50 is plausibly caused by foreground absorption. To do so, we used the 3D dust maps from Bayestar19 (see Green et al. 2019, for detailed description). We are looking for the intervening absorption between W50 and us, i.e., physically unrelated layers of the absorbing material. Given the adopted distance to SS 433 of 5 kpc, we integrated extinction () only up to a distance of 3 kpc from the Sun. The resulting extinction map is superposed on the X-ray image in Fig. 9. A prominent wide dust lane is seen in this image that crosses the central part of the W50 nebula diagonally from NE to SW. Typical values of the extinction in this region are , which translate into equivalent hydrogen column densities . Therefore, soft X-ray emission from the central part of W50 can be heavily absorbed (several orders of magnitude at 1 keV). While some geometric correspondences between the X-ray and extinction maps look tantalizing, we conclude that most probably this is a by-chance projection. Therefore, the lack of observed soft X-ray emission from the quasi-circular part of the nebula does not place significant constraints on the properties of W50. The soft X-ray emission might be there, but we can not see it.

### v2

@@
 X-ray panorama of the SS 433/W50 complex by SRG/eROSITA
-Galactic microquasar SS 433 and the radio nebula W50 surrounding it present a prototypical example of a hyper-Eddington binary system shaping its ambient interstellar medium via energetic outflows.
+The Galactic microquasar SS 433 and the radio nebula W50 surrounding it present a prototypical example of a hyper-Eddington binary system shaping its ambient interstellar medium via energetic outflows.
 In this paper, we present X-ray observations of the SS 433/W50 complex by the eROSITA telescope onboard the SRG space observatory.
 These data provide images of the entire nebula characterized by a very large dynamic range and allow spectral analysis of the diffuse X-ray emission.
-In particular, these data illustrate a close connection between the thermal and non-thermal components of W50 on scales ranging from sub-parsecs, represented by narrow X-ray bright filaments, to the entire extent of the nebula.
-These data also allow us to fully characterize a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts, but are prominent in very high energy gamma-ray emission.
+In particular, these data illustrate a close connection between the thermal and non-thermal components of W50 on scales ranging from sub-parsecs, represented by narrow X-ray bright filaments, to the entire extent of of the nebula.
+These data also allowed us to fully characterize a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts but are prominent in very-high-energy gamma-ray emission.
 The resulting multifaceted picture of the interaction between energetic outflows and the surrounding medium paves the way for future focused multiwavelength observations and dedicated numerical simulations.
-Key Words.:
-ISM: jets and outflows – X-rays: binaries – plasmas – acceleration of particles1 Introduction
-SS 433 is a unique close stellar binary system in our Galaxy (located at and kpc) composed of a relativistic compact object (a black hole or a neutron star) which accretes matter from its companion at sustainably hyper-Eddington rate (see Fabrika, 2004; Cherepashchuk et al., 2025, for reviews).
-Having been first listed in the catalog of point-like H emitters (Stephenson & Sanduleak, 1977), it quickly became one of the most studied astrophysical objects thanks to its very peculiar optical spectrum (Margon et al., 1979) featuring pairs of narrow (with half-opening angle ) blue- and red-shifted lines (with respect to their rest frame positions) with the shifts varying periodically in time (see Margon, 1984, for a review of early observational results).
-A kinematic model associates this emission with a pair of narrow relativistic jets having constant bulk velocity of a quarter of the speed of light and changing direction in a regular precession manner with a period days (Abell & Margon, 1979; Fabian & Rees, 1979).
-This period is much longer than the period of orbital modulation ( days) and is interpreted as due to precession of a thick accretion disk with the amplitude of 21∘ and the mean axis inclined by the angle to the line of sight (e.g., Katz, 1980; Sarazin et al., 1980; Whitmire & Matese, 1980).
-Subsequent detection of extended and variable radio emission (Spencer, 1979) perfectly described by the precessing jets pattern fully confirmed the picture of twin relativistic outflows (Hjellming & Johnston, 1981a, b; Blundell & Bowler, 2004).
-Moreover, the proper motion of the individual radio-emitting blobs allowed one to constrain the distance to the source with a remarkable 10% accuracy, being kpc (Hjellming & Johnston, 1981a; Blundell & Bowler, 2004; Marshall et al., 2013).
-Despite indications of both quasi-regular (nutation, Katz et al., 1982) and sporadic (jittering, Kubota et al., 2010) deviations from the predictions of this simple model as well as presence of (relatively short) periods of time when the Doppler shifts of optical emission lines change significantly (e.g., Blundell et al., 2011; Medvedev et al., 2022), all the derived parameters of the model stay remarkably stable over more than 40 years of continuous monitoring observations (Cherepashchuk et al., 2018).
-This stability likely reflects extreme robustness of the mass transfer occurring in the hyper-Eddington regime (van den Heuvel, 1981; van den Heuvel et al., 2017) with the estimated accretion rate (e.g., Shklovskii, 1981; Cherepashchuk, 1981; Leibowitz, 1984; Fuchs et al., 2006), where is the mass of the compact object and is the Solar mass and is the critical Eddington mass accretion rate.
-The compact object is most likely a black hole with (Cherepashchuk et al., 2023), although the possibility of it being a neutron star () has also been discussed (e.g.
-Goranskij, 2011; Medvedev et al., 2018).
-Since X-ray emission coming from the innermost parts of the hyper-Eddington accretion disk is expected to be strongly beamed in the axial direction (Shakura & Sunyaev, 1973; Begelman et al., 2006; Poutanen et al., 2007), we are not able to observe it directly, given the high inclination angle of the system.
-Indeed, although X-ray emission from this source has also been detected early on (Seward et al., 1976; Forman et al., 1978), its luminosity is relatively low compared to the Eddington luminosity (Grindlay et al., 1984).
-As a result, SS 433 has long been considered as a beamed away prototype of Ultra Luminous X-ray sources (ULXs, e.g., Kaaret et al., 2017, for a review) observed in normal star-forming galaxies (Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007), although possible indications of the ‘hidden’ central emission are yet to be confirmed (Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016, 2019; Middleton et al., 2021; Fogantini et al., 2023).
-Instead, what dominates X-ray emission of the system is again thermal emission from a twin pair of the baryonic jets having almost identical orientation and bulk velocity to the optically emitting jets (Watson et al., 1986).
-Numerous pairs of blue- and red-shifted narrow emission lines of highly ionized heavy elements (primarily, silicon, sulfur, iron and nickel) allow one to build a multi-temperature model of the ballistically expanding and cooling flow (Watson et al., 1986; Kotani et al., 1996; Marshall et al., 2002; Brinkmann et al., 2005; Medvedev & Fabrika, 2010; Marshall et al., 2013; Khabibullin et al., 2016) which first becomes visible when its temperature is above 20 keV and then cools down below 1 keV and likely fragments later on due to thermal instability (Brinkmann et al., 1988).
-Detailed spectroscopic modeling of this emission (e.g.
-Khabibullin et al., 2016; Medvedev et al., 2019) in combination with the eclipsing observations of the jets by the companion star (Stewart et al., 1987; Kawai et al., 1989; Filippova et al., 2006; Lopez et al., 2006; Atapin & Fabrika, 2016) enable very robust determination of the jets’ physical properties (so-called jets ‘biometry’, consisting of their size, opening angle, velocity, temperature, and the mass flux).
-Thanks to this we know very robustly the kinetic luminosity of these outflows, which exceeds (e.g., Medvedev et al., 2019), confirming the necessity of a hyper-Eddington ‘engine’ releasing large amount of energy in the form of relativistic outflows, consistently with the early theoretical (Shakura & Sunyaev, 1973) and modern numerical (e.g., Ohsuga & Mineshige, 2011; Jiang et al., 2019; Yoshioka et al., 2022; Fragile et al., 2025) predictions.
+1 Introduction
+SS 433 is a unique close stellar binary system in our Galaxy (located at and kpc) composed of a relativistic compact object (a black hole or a neutron star), which sustainably accretes matter from its companion at a hyper-Eddington rate (see Fabrika, 2004; Cherepashchuk et al., 2025, for reviews).
+Having first been listed in the catalog of point-like H emitters (Stephenson and Sanduleak, 1977), it quickly became one of the most studied astrophysical objects thanks to its very peculiar optical spectrum (Margon et al., 1979) featuring pairs of narrow (with a half-opening angle of ) blue- and redshifted lines (with respect to their rest-frame positions), with the shifts varying periodically in time (for a review of early observational results see Margon, 1984).
+A kinematic model associates this emission with a pair of narrow relativistic jets with a constant bulk velocity of a quarter of the speed of light (c) that change direction in a regular precession manner with a period of days (Abell and Margon, 1979; Fabian and Rees, 1979).
+This period is much longer than the period of orbital modulation ( days) and is interpreted as being due to the precession of a thick accretion disk with an amplitude of 21∘ and mean axis inclined by the angle to the line of sight (e.g., Katz, 1980; Sarazin et al., 1980; Whitmire and Matese, 1980).
+Subsequent detection of extended and variable radio emission (Spencer, 1979) perfectly described by the precessing jets pattern fully confirmed the picture of twin relativistic outflows (Hjellming and Johnston, 1981a, b; Blundell and Bowler, 2004).
+Moreover, the proper motion of the individual radio-emitting blobs allowed one to constrain the distance to the source with a remarkable 10% accuracy; i.e., kpc (Hjellming and Johnston, 1981a; Blundell and Bowler, 2004; Marshall et al., 2013).
+Despite indications of both quasi-regular (nutation, Katz et al., 1982) and sporadic (jittering, Kubota et al., 2010) deviations from the predictions of this simple model as well as the presence of (relatively short) periods of time when the Doppler shifts of optical emission lines change significantly (e.g., Blundell et al., 2011; Medvedev et al., 2022), all the derived parameters of the model stay remarkably stable over more than 40 years of continuous monitoring observations (Cherepashchuk et al., 2018).
+This stability likely reflects extreme robustness of the mass transfer occurring in the hyper-Eddington regime (van den Heuvel, 1981; van den Heuvel et al., 2017) with the estimated accretion rate, (e.g., Shklovskii, 1981; Cherepashchuk, 1981; Leibowitz, 1984; Fuchs et al., 2006), where is the mass of the compact object, is the solar mass, and is the critical Eddington mass-accretion rate.
+The compact object is most likely a black hole with (Cherepashchuk et al., 2023), although the possibility of it being a neutron star () has also been discussed (e.g., Goranskij, 2011; Medvedev et al., 2018).
+Since X-ray emission coming from the innermost parts of the hyper-Eddington accretion disk is expected to be strongly beamed in the axial direction (Shakura and Sunyaev, 1973; Begelman et al., 2006; Poutanen et al., 2007), we are not able to observe it directly, given the high inclination angle of the system.
+Although X-ray emission from this source was detected early on (Seward et al., 1976; Forman et al., 1978), it was also found (e.g., Grindlay et al., 1984) that its apparent luminosity, , is very low compared to the Eddington luminosity, erg s-1, of an compact object.
+As a result, SS 433 has long been considered as a beamed away prototype of ultra-luminous X-ray sources (ULXs; for a review see Kaaret et al., 2017) observed in normal star-forming galaxies (Fabrika and Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007), although possible indications of the ”hidden” central emission are yet to be confirmed (Medvedev and Fabrika, 2010; Khabibullin and Sazonov, 2016, 2019; Middleton et al., 2021; Fogantini et al., 2023).
+Instead, what dominates the X-ray emission of the system is again thermal emission from a twin pair of the baryonic jets having almost identical orientation and bulk velocity to the optically emitting jets (Watson et al., 1986).
+Numerous pairs of blue- and redshifted narrow emission lines of highly ionized heavy elements (primarily silicon, sulfur, iron, and nickel) allow one to build a multi-temperature model of the ballistically expanding and cooling flow (Watson et al., 1986; Kotani et al., 1996; Marshall et al., 2002; Brinkmann et al., 2005; Medvedev and Fabrika, 2010; Marshall et al., 2013; Khabibullin et al., 2016), which first becomes visible when its temperature is above 20 kiloelectronvolt (keV), then cools down below 1 keV, and likely fragments later on due to thermal instability (Brinkmann et al., 1988).
+Detailed spectroscopic modeling of this emission (e.g., Khabibullin et al., 2016; Medvedev et al., 2019) in combination with the eclipsing observations of the jets by the companion star (Stewart et al., 1987; Kawai et al., 1989; Filippova et al., 2006; Lopez et al., 2006; Atapin and Fabrika, 2016) enable very robust determination of the jets’ physical properties (their so-called ”biometry”, consisting of their size, opening angle, velocity, temperature, and the mass flux).
+Thanks to this we know the kinetic luminosity of these outflows with certainty; it exceeds (e.g., Medvedev et al., 2019), confirming the necessity of a hyper-Eddington ”engine” releasing large amount of energy in the form of relativistic outflows, consistently with the early theoretical (Shakura and Sunyaev, 1973) and modern numerical (e.g., Ohsuga and Mineshige, 2011; Jiang et al., 2019; Yoshioka et al., 2022; Fragile et al., 2025) predictions.
 In contrast to the radiation freely escaping the system, the kinetic energy of the launched outflows cannot be easily transported away from the source and has to be deposited into the surrounding interstellar medium.
-The estimated total amount of this energy exceeds erg for a life-time of the system at the level of 30,000 yr (e.g., van den Heuvel et al., 1980).
-This energy is at least on par with the energy of the initial supernova explosion that should have accompanied the birth of the compact object, meaning that a quite large region pc in its vicinity can be shaped by this later activity (e.g., Begelman et al., 1980; Zealey et al., 1980; van den Heuvel, 1981; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Churazov et al., 2020, 2024).
-Indeed, SS 433 is known to be surrounded by the famous giant ( pc) radio nebula W50 (Westerhout, 1958; Holden & Caswell, 1969; Geldzahler et al., 1980; Elston & Baum, 1987; Dubner et al., 1998; Goodall et al., 2011b; Broderick et al., 2018; Sakemi et al., 2021), which has a suggestive elongated morphology, with the elongation axis being co-aligned with the symmetry axis of the jets precession cone.
-The asymmetry of the lobes, or “ears”, of the nebula can be readily explained by the noticeable stratification of the surrounding medium, given that the system is located at the height pc below the Galactic plane and has the size comparable to the scale height of the exponential cold gas disk (e.g., Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a).
-The interior of the nebula is also known to shine in soft thermal X-rays (Brinkmann et al., 1996, 2007; Safi-Harb & Ögelman, 1997; Chi et al., 2024), hard non-thermal (Seward et al., 1980; Watson et al., 1983; Yamauchi et al., 1994; Safi-Harb & Petre, 1999; Namiki et al., 2000; Safi-Harb et al., 2022; Kaaret et al., 2024) and filamentary H emission (Kirshner & Chevalier, 1980; Shuder et al., 1980; Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017; Rosado et al., 2021), indicative of the high velocity shocks propagating through a relatively dense interstellar medium.
-The nature of the extended X-ray jets (EXJ) remains more puzzling, although their morphology, spectra, and polarization all point to the synchrotron origin of the emission (e.g., Safi-Harb et al., 2022; Kaaret et al., 2024).
-Most recently, Very High Energy (VHE) emission from these structures has been detected and mapped (e.g., Abeysekara et al., 2018; Alfaro et al., 2024; H.
+The estimated total amount of this energy exceeds erg for a lifetime of the system at the level of 30,000 yr (e.g., van den Heuvel et al., 1980).
+This energy is at least on par with the energy of the initial supernova explosion that accompanied the birth of the compact object, meaning that quite a large region of pc in its vicinity can be shaped by this later activity (e.g., Begelman et al., 1980; Zealey et al., 1980; van den Heuvel, 1981; Velázquez and Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Churazov et al., 2020, 2024).
+Indeed, SS 433 is known to be surrounded by the famous giant ( pc) radio nebula W50 (Westerhout, 1958; Holden and Caswell, 1969; Geldzahler et al., 1980; Elston and Baum, 1987; Dubner et al., 1998; Goodall et al., 2011b; Broderick et al., 2018; Sakemi et al., 2021); the latter has a suggestive, elongated morphology, with the elongation axis being co-aligned with the symmetry axis of the jets’ precession cone.
+The asymmetry of the lobes, or “ears”, of the nebula can be readily explained by the noticeable stratification of the surrounding medium, given that the system is located at a height of pc below the Galactic plane and has the size comparable to the scale height of the exponential cold gas disk (e.g., Velázquez and Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a).
+The interior of the nebula is also known to shine in soft thermal X-rays (Brinkmann et al., 1996, 2007; Safi-Harb and Ögelman, 1997; Chi et al., 2024), hard nonthermal X-rays (Seward et al., 1980; Watson et al., 1983; Yamauchi et al., 1994; Safi-Harb and Petre, 1999; Namiki et al., 2000; Safi-Harb et al., 2022; Kaaret et al., 2024), and filamentary H emission (Kirshner and Chevalier, 1980; Shuder et al., 1980; Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017; Rosado et al., 2021), which is indicative of the high-velocity shocks propagating through a relatively dense interstellar medium.
+The nature of the extended X-ray jets (EXJs) remains more puzzling, although their morphology, spectra, and polarization all point to the synchrotron origin of the emission (e.g., Safi-Harb et al., 2022; Kaaret et al., 2024).
+Most recently, very-high-energy (VHE) emission from these structures has been detected and mapped (e.g., Abeysekara et al., 2018; Alfaro et al., 2024; H.
 E.
@@
 S.
-Collaboration et al., 2024; LHAASO Collaboration, 2024), demonstrating that indeed ultra-relativistic electrons with energy above 100 TeV are accelerated in the system (e.g, Kimura et al., 2020), probably as a consequence of internal shocks and recollimation in the energetic trans-relativistic outflows (Bykov et al., 2025).
-The overall production of cosmic rays by such systems (super-Eddington accretors with powerful outflows) might be of importance for the total CR budget and local variations in its spectrum from place to place (e.g., Heinz & Sunyaev, 2002), in particular in the PeV regime (e.g., Peretti et al., 2025; Wang et al., 2025; Bykov et al., 2025).
+Collaboration et al., 2024; LHAASO Collaboration, 2025), demonstrating that ultra-relativistic electrons with energy above 100 teraelectronvolt (TeV) are accelerated in the system (e.g., Kimura et al., 2020), probably as a consequence of internal shocks and recollimation in the energetic trans-relativistic outflows (Bykov et al., 2025).
+The overall production of cosmic rays by such systems (super-Eddington accretors with powerful outflows) might be of importance for the total CR budget and local variations in its spectrum from place to place (e.g., Heinz and Sunyaev, 2002), particularly in the petaelectronvolt (PeV) regime (e.g., Peretti et al., 2025; Wang et al., 2025; Bykov et al., 2025).
 Although the paradigm of “a super-Eddington X-ray binary powering a nebula with energetic outflows” appears to be well established and consistent with the most salient observational properties of the SS 433/W50 complex, the details of almost all aspects of this interaction remain unclear.
-Indeed, all the phenomena directly associated with the central source can be trace only up to a distance of mere 0.1 pc from it, where the last signatures of the corkscrew pattern of the precessing jets are observed in radio and X-rays (Migliari et al., 2005; Miller-Jones et al., 2008; Khabibullin & Sazonov, 2017; Sakai et al., 2025).
-A model invoking steady deceleration and disruption of the jets can, in principle, account for such behavior (Panferov, 2014, 2017), but then the re-appearance of the “jets” with similar to the injection velocities tens of pc away from the central source might be problematic.
-Morphology of the soft X-ray emission, best mapped by the ROSAT mosaic observations (Brinkmann et al., 1996, 2005), is also quite puzzling, with large global brightness variations and thin filamentary structures, not very commonly observed in similar situations of supernova remnants (which typically feature shell-, plerion- or mixed types of morphology).
-Deep pointed observations with XMM-Newton, Chandra, Suzaku, and NuSTAR were capable of covering certain parts of the nebula (e.g., Safi-Harb et al., 2022; Chi et al., 2024; Tsuji et al., 2025), but a uniform and sensitive coverage of the entire nebula with adequate spectroscopic capabilities is missing.
-Here we present results of X-ray observations of the SS 433/W50 complex by the eROSITA telescope (Predehl et al., 2021) onboard the SRG space observatory (Sunyaev et al., 2021), which were performed during its Performance Verification (PV) phase.
-The obtained mosaic provides us with X-ray images of the entire extent of W50 with unprecedented dynamic range, bringing X-ray data on par with the radio and other multi-wavelength data, and, in addition, delivers exquisite spectral information for the rich variety of diffuse structures.
-These data, in particular, demonstrate a close connection between the thermal and non-thermal components of W50 on scales ranging from subparsec, represented by narrow X-ray and radio bright filaments, to the entire extent of the nebula.
-These data also enable full characterization of a pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system, which lack radio counterparts.
-The structure of the paper is following: we describe the observations and the obtained data in Sect.
-2; results of the broad-band and spectrally-resolved imaging are presented in Sect.
+Indeed, all the phenomena directly associated with the central source can only be traced up to a distance of 0.1 pc from it, where the last signatures of the corkscrew pattern of the precessing jets are observed in radio and X-rays (Migliari et al., 2005; Miller-Jones et al., 2008; Khabibullin and Sazonov, 2017; Sakai et al., 2025).
+A model invoking steady deceleration and disruption of the jets can, in principle, account for such behavior (Panferov, 2014, 2017), but then the reappearance of the “jets” tens of parsec away from the central source with velocities similar to those at the injection might be problematic.
+Morphology of the soft X-ray emission, best mapped by the ROSAT mosaic observations (Brinkmann et al., 1996), is also quite puzzling, with large global brightness variations and thin filamentary structures, which are not commonly observed in similar situations of supernova remnants (SNRs, which typically feature shell-, plerion- or mixed types of morphology).
+Deep pointed observations with XMM-Newton, Chandra, Suzaku, and NuSTAR were capable of covering certain parts of the nebula (e.g., Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024; Tsuji et al., 2025), but a uniform and sensitive coverage of the entire nebula with adequate spectroscopic capabilities is missing.
+Here, we present results of X-ray observations of the SS 433/W50 complex by the eROSITA telescope (Predehl et al., 2021) on board the SRG space observatory (Sunyaev et al., 2021), which were performed during its performance-verification (PV) phase.
+The obtained mosaic provides us with X-ray images of the entire extent of W50 with unprecedented dynamic range, bringing X-ray data on par with the radio and other multiwavelength data, and, in addition, delivers exquisite spectral information for the rich variety of diffuse structures.
+These data demonstrate a close connection between the thermal and nonthermal components of W50 on scales ranging from subparsec (thickness of narrow X-ray and radio bright filaments), to the entire pc extent of the nebula.
+These data also enable the characterization of the pair of nearly symmetric, sharp-edged, elongated structures aligned with the orbital axis of the binary system (”extended X-ray jets”).
+The structure of the paper is as follows.
+We describe the observations and the obtained data in Sect.
+2; results of the broad-band and spectrally resolved imaging are presented in Sect.
 3 and Sect.
-4, respectively; spectroscopic analysis of the emission from several selected regions follows in Sect.
-5; the obtained global picture is discussed in Sect.
+4, respectively; a spectroscopic analysis of the emission from several selected regions follows in Sect.
+5; and the obtained global picture is discussed in Sect.
 6 and summarized in Sect.
@@
 2 Data
-Since W50 spans approximately degrees on the sky, corresponding to the physical size of at the distance of 5 kpc111The most recent parallax measurement for SS 433 in Gaia DR3 corresponds to the significantly larger distance of kpc (Gaia DR3 4293406612283985024, Gaia Collaboration et al., 2023), we caution of using this value given large residuals of the astrometric solution.
-The previous parallax measurement published in Gaia DR2 corresponded to a much smaller distance of kpc (Gaia DR2 4293406612283985024, Arnason et al., 2021)., the observing campaign consisted of 9 individual telescope pointings (each covering a circular area of in diameter).
-These pointings were arranged in a mosaic to result in a nearly uniform coverage of the entire nebula and, in addition, the highest sensitivity in the most interesting regions.
-The observing dates, from October 5th to October 8th and then from October 19th to October 20th, 2019, were chosen to coincide with the two consecutive orbital eclipses of the central source SS 433 by its normal companion star (according to orbital ephemerids by Eikenberry et al., 2001).
-During eclipses, the apparent X-ray flux of SS 433 goes down by a factor of several (e.g., Kawai et al., 1989; Lopez et al., 2006; Medvedev et al., 2019), minimizing possible adverse effects of contamination and pile-up due to the presence of a very bright object in the field-of-view, even though relative suppression of the soft X-ray lines being rather minor (Marshall et al., 2013).
-The typical effective (i.e., equivalent to the nominal eROSITA on-axis performance) exposure time across the map amounts to 30-50 ks, leading to a factor of deeper images compared to the previously available full-extent maps of W50 by observatory in the 0.5-2 keV band (Brinkmann et al., 1996) and Einstein observatory in the 2-4 keV band (Watson et al., 1983).
-In addition to that, the eROSITA data have much better spatial ( arcsec on the telescope axis and averaged across the field-of-view) and spectral ( eV at 1 keV) resolution (Predehl et al., 2021).
-A comparable quality of the data is provided by multiple Chandra and XMM-Newton observations, but their coverage of the whole W50 nebula is non-uniform and still incomplete after more than 20 years of observations (Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024).
-On the other hand, the SRG/eROSITA data are complementary to these observations, as well as to the imaging observations in harder X-rays with NuSTAR, which offer a unique view of non-thermally emitting populations.
+Since W50 spans approximately degrees on the sky, corresponding to a physical size of at a distance of 5 kpc111The most recent parallax measurement for SS 433 in Gaia DR3 corresponds to much larger distance of kpc (Gaia DR3 4293406612283985024, Gaia Collaboration et al., 2023); we advise caution if using this value given the large residuals of the astrometric solution.
+The previous parallax measurement from Gaia DR2 data gave a much smaller distance of kpc (Arnason et al., 2021)., the observing campaign consisted of nine individual telescope pointings (each covering a circular area of in diameter).
+These pointings were arranged in a mosaic to give a nearly uniform coverage of the entire nebula and, in addition, the highest sensitivity in the most interesting regions.
+The observing dates, from October 5 to October 8 and then from October 19 to October 20, 2019, were chosen to coincide with the two consecutive orbital eclipses of the central source SS 433 by its normal companion star (according to orbital ephemerids by Eikenberry et al., 2001).
+During eclipses, the apparent X-ray flux of SS 433 goes down by a factor of several (e.g., Kawai et al., 1989; Lopez et al., 2006; Medvedev et al., 2019), minimizing possible adverse effects of contamination and pile-up due to the presence of a very bright object in the field of view, even though the relative suppression of the soft X-ray lines was rather minor (Marshall et al., 2013).
+The typical effective (i.e., equivalent to the nominal eROSITA on-axis performance) exposure time across the map amounts to 30-50 ks, leading to images a factor of ten deeper compared to the previously available full-extent maps of W50 by the observatory in the 0.5-2 keV band (Brinkmann et al., 1996) and Einstein observatory in the 2-4 keV band (Watson et al., 1983).
+In addition to that, the eROSITA data have much better spatial ( on the telescope axis and averaged across the field of view) and spectral ( eV at 1 keV) resolution (Predehl et al., 2021).
+A comparable quality of the data is provided by multiple Chandra and XMM-Newton observations, but their coverage of the whole W50 nebula is nonuniform and still incomplete after more than 20 years of observations (Brinkmann et al., 2005; Safi-Harb et al., 2022; Chi et al., 2024).
+On the other hand, the SRG/eROSITA data are complementary to these observations, as well as to the imaging observations in harder X-rays with NuSTAR, which offer a unique view of nonthermally emitting populations.
 In this paper, we focus on the SRG/eROSITA X-ray data alone and leave combined analyses for future studies.
 We also supplement the PV data with the data obtained in the course of the SRG/eROSITA All-Sky Survey (Predehl et al., 2021; Sunyaev et al., 2021).
-Although the survey data are much shallower in sensitivity (the typical exposure time per point is s after combining the data of four scans), they provide uniform coverage of the entire area surrounding the nebula.
-For them, the standard data reduction, background estimation and exposure- and vignetting corrections were applied in the same way as was done for mapping diffuse emission from nearby galaxy clusters (Churazov et al., 2021a, 2023), supernova remnants (Churazov et al., 2021b; Khabibullin et al., 2023, 2024), and the Galactic diffuse emission (Predehl et al., 2020; Khabibullin et al., 2022; Churazov et al., 2024) on even larger scales.
-Here we also go one step further and combine the PV and all-sky data in a seamless fashion, similarly as it has been done for Chandra and SRG/eROSITA observations of the Perseus cluster (Churazov et al., 2025).
-A comparison of the degrees field in radio and X-ray bands is shown in Fig.
+Although the survey data are much shallower (the typical exposure time per point is s after combining the data of four scans), they provide uniform coverage of the entire area surrounding the nebula.
+For them, the standard data reduction, background estimation, and exposure and vignetting corrections were applied (similarly to mapping diffuse emission from nearby galaxy clusters (Churazov et al., 2021a, 2023), supernova remnants (Churazov et al., 2021b; Khabibullin et al., 2023, 2024), and the Galactic diffuse emission (Predehl et al., 2020; Khabibullin et al., 2022; Churazov et al., 2024) on even larger scales).
+Here, we also went one step further and combined the PV and all-sky data in a seamless fashion, similarly as it has been done for Chandra, XMM-Newton and SRG/eROSITA observations of the Perseus cluster (Churazov et al., 2025).
+A comparison of the degree fields in radio and X-ray bands is shown in Fig.
 1.
-Although in the radio image, the outer boundary of the W50 nebula appears to be very prominent, the X-ray image is instead dominated by the very bright central object (SS 433) and elongated structures known as “extended X-ray jets” (EXJ) located well inside the radio boundary of W50 (Seward et al., 1980; Watson et al., 1983; Brinkmann et al., 1996).
-At a distance of 5 kpc, these images span 350 pc and range from 17 pc to almost 370 pc in distance from the Galactic plane.
+Although in the radio image the outer boundary of the W50 nebula appears to be very prominent, the X-ray image is instead dominated by the very bright central object (SS 433) and elongated structures known as extended X-ray jets (EXJs) located well inside the radio boundary of W50 (Seward et al., 1980; Watson et al., 1983; Brinkmann et al., 1996).
+At a distance of 5 kpc, these images span 350 pc and range from 17 pc to almost 370 pc in terms of distance from the Galactic plane.
 Given that the scale heights for the cold and warm components of the interstellar medium are tens and hundreds of pc, respectively, one could naturally expect signatures of the environment stratification (e.g., Goodall et al., 2011a).
-This is indeed revealed by the asymmetry of the W50 lobes, with the one directed towards the Galactic plane (the Western Lobe in equatorial coordinates) being more compact than the opposite one (the Eastern lobe).
+This is indeed revealed by the asymmetry of the W50 lobes, with the one directed toward the Galactic plane (the western lobe in equatorial coordinates) being more compact than the opposite one (the eastern lobe).
 On larger scales, clear variations of the background and foreground diffuse X-ray emission are seen.
-Inspection of 3D reddening maps (e.g., Green et al., 2019) shows a clear anti-correlation of the reddening and soft X-ray surface brightness (see Appendix C), suggesting that absorption of X-rays is the main reason for these variations.
+An inspection of 3D reddening maps (e.g., Green et al., 2019) showed a clear anticorrelation of the reddening and soft X-ray surface brightness (see Appendix C), suggesting that absorption of X-rays is the main reason for these variations.
 Morphologically similar structures appear in the 3D reddening maps at a distance of 2-3 kpc, i.e., in the foreground to W50/SS 433.
@@
 2, with the contours of radio emission at 4.85 GHz (Gregory et al., 1996) overlaid in white.
-The color coding reflects the measured X-ray surface brightness on a square-root scale spanning two orders of magnitude, highlighting the wide dynamic range of the diffuse X-ray emission captured by eROSITA.
-In addition to SS 433 itself (the extremely bright source in the center) and a multitude of point sources, the majority of which are nearby foreground stars or background AGN, the picture shows bright diffuse emission with rich internal structure.
+The color-coding reflects the measured X-ray surface brightness on a square-root scale spanning two orders of magnitude, highlighting the wide dynamic range of the diffuse X-ray emission captured by eROSITA.
+In addition to SS 433 itself (the extremely bright source in the center) and a multitude of point sources –the majority of which are nearby foreground stars or background active galactic nuclei (AGNs) – the picture shows bright diffuse emission with rich internal structure.
 The X-ray picture can be broadly decomposed into three components.
-Firstly, there is a low surface brightness emission filling the entire projected extent of the nebula.
+First, there is a low-surface-brightness emission filling the entire projected extent of the nebula.
 It is brighter in the eastern and western lobes of W50 (left and right, respectively, in Fig.
@@
 Remarkably, the central part of W50 (beyond the region contaminated by the PSF wings from SS 433) appears even fainter in X-rays, suggesting a shell-like geometry of the X-ray-emitting gas.
-Secondly, one can notice the presence of very narrow X-ray filaments, particularly inside and close to the boundaries of the Eastern and Western lobes.
+Second, one can notice the presence of very narrow X-ray filaments, particularly inside and close to the boundaries of the eastern and western lobes.
 Their apparent width is , while their length is .
-Given the angular resolution of eROSITA (better than , i.e., at 5 kpc), these filaments are fully resolved in both directions.
+Given the angular resolution of eROSITA (better than ; i.e., at 5 kpc), these filaments are fully resolved in both directions.
 The X-ray filaments form a complicated network with an overall structure very similar to the filaments of the cold and dense gas observed via H+[N II] emission at different locations within the nebula (Boumis et al., 2007; Farnes et al., 2017).
 The largest X-ray filaments have a flat transverse profile of the surface brightness and are an order of magnitude brighter than the surrounding medium.
-Such filaments of the dense X-ray emitting gas might arise as instabilities in the sheared flows, caustics of the strong shock waves, or due to heating of the preexisting cold gas filaments by hot gas and relativistic particles inside the lobes.
-The final and perhaps the most visual feature of the image is a pair of nearly symmetrical, sharp-edged “extended X-ray jets” directed away from the central source.
-Their axis is not only aligned with the elongation of the radio nebula and location of the X-ray bright lobes, but also coincides with the projection of the precession axis of SS 433’s famous trans-relativistic baryonic jets.
-The physical nature of these structures, first found by the Einstein observatory (Watson et al., 1983) and then observed with all major X-ray telescopes (Brinkmann et al., 1996; Safi-Harb & Ögelman, 1997; Namiki et al., 2000; Brinkmann et al., 2007; Safi-Harb & Petre, 1999), including the most recent polarimteric observations with IXPE (Kaaret et al., 2024), remains a matter of debate (e.g.
-Churazov et al., 2024).
-The SRG/eROSITA data clearly show a very rapid onset of emission at the innermost boundary of EXJs on both sides of the nebula (also demonstrated recently by Tsuji et al., 2025, using Chandra data).
-4 Spectrally-decomposed imaging
-We take advantage of the unique possibility provided by /eROSITA to probe the spectral characteristics of diffuse X-ray emission throughout the full extent of the nebula for the first time.
-To produce a spectrally-decomposed image, we divided the full spectral band presented earlier into three sub-bands: soft (0.5-1 keV), medium (1-2 keV), and hard (2-4 keV) bands, respectively.
-A composite RGB image showing the surface brightness of the emission (now on a linear scale) in these bands in red, green, and blue, respectively, is presented in Figure 3.
-The chosen spectral decomposition immediately reveals the presence of two distinct components in the observed diffuse emission: large regions of red-orange color where emission below 1 keV dominates, and green-cyan structures where prominent emission above 1 keV is present (including the bright central source SS 433).
-Compact sources also form two groups: the softer (“red-yellow”) ones are predominantly foreground stars, while harder (“blueish”) sources are mostly background AGNs, which shine through the gas and dust in the disk of the Milky Way.
-Zoom-in images of the lobes (rotated and put side by side to help the comparison) are shown in Fig.
+Such filaments of the dense X-ray-emitting gas might arise as instabilities in the sheared flows, caustics of the strong shock waves, or due to heating of the preexisting cold-gas filaments by hot gas and relativistic particles inside the lobes.
+The final and perhaps the most visual feature of the image is a pair of nearly symmetrical, sharp-edged EXJs directed away from the central source.
+Their axis is not only aligned with the elongation of the radio nebula and location of the X-ray bright lobes, it also coincides with the projection of the precession axis of SS 433’s famous trans-relativistic baryonic jets.
+The physical nature of these structures, first found by the Einstein observatory (Watson et al., 1983) and then observed with all major X-ray telescopes (Brinkmann et al., 1996; Safi-Harb and Ögelman, 1997; Namiki et al., 2000; Brinkmann et al., 2007; Safi-Harb and Petre, 1999), including the most recent polarimteric observations with IXPE (Kaaret et al., 2024), remains a matter of debate (e.g., Churazov et al., 2024).
+The /eROSITA data clearly show a very rapid onset of emission at the innermost boundary of EXJs on both sides of the nebula (also demonstrated recently using Chandra data by Tsuji et al., 2025).
+4 Spectrally decomposed imaging
+We took advantage of the unique possibility provided by /eROSITA to probe the spectral characteristics of diffuse X-ray emission throughout the full extent of the nebula for the first time.
+To produce a spectrally decomposed image, we divided the full spectral band presented earlier into three subbands: soft (0.5-1 keV), medium (1-2 keV), and hard (2-4 keV) bands, respectively.
+A composite pseudo-color (red-green-blue) image showing the surface brightness of the emission (now on a linear scale) in these bands in red, green, and blue, respectively, is presented in Figure 3.
+The chosen spectral decomposition immediately reveals the presence of two distinct components in the observed diffuse emission: large regions in red-orange where emission below 1 keV dominates, and green-cyan structures where prominent emission above 1 keV is present (including the bright central source SS 433).
+Compact sources also form two groups: the softer (red-yellow) ones are predominantly foreground stars, while harder (blueish) sources are mostly background AGNs, which shine through the gas and dust in the disk of the Milky Way.
+Zoomed-in images of the lobes (rotated and put side by side to aid comparison) are shown in Fig.
 4 and demonstrate fine structure of the soft diffuse emission, contrasting with relatively smooth and well-confined hard emission.
@@
 The overlaid distance and angle rulers highlight the characteristic dimensions and morphology of the prominent structures.
-Namely, one can see that the “Extended X-ray Jets” start pc away from the central source and demonstrate remarkable symmetry, with the Eastern one starting only a couple of parsecs further from the central source.
-Both EXJs, characterized by hard X-ray spectra, terminate around pc.
+Namely, one can see that the EXJs’ start pc away from the central source and demonstrate remarkable symmetry, with the eastern one starting only a couple of parsecs further from the central source.
+Both EXJs, characterized by hard X-ray spectra, terminate at around pc.
 This symmetry contrasts with the more asymmetric and irregular soft emission of the lobes.
@@
 Indeed, while the eastern lobe extends up to , the western one terminates at pc.
-The inclination of the relativistic jet’s precession axis to the line of sight is (the relativistic jet on the eastern side is on average approaching us), so this difference cannot be explained by geometrical effects only.
+The inclination of the relativistic jet’s precession axis to the line of sight is (the relativistic jet on the eastern side is on average approaching us), so this difference cannot be explained by geometrical effects alone.
 However, it broadly agrees with the assumption that the energy flow from the central source interacts with the denser ambient medium in the western part, reflecting the gas density gradient perpendicular to the equatorial plane of the Galactic disk (e.g., Goodall et al., 2011a).
-Since the difference between the lobes’ dimensions amounts to a sizeable fraction of them, one can infer the exponential scale-height of the ambient gas density profile pc, consistent with the scale-height of the Galactic neutral gas disk.
-The X-ray emission from the Northern and Southern boundaries of the nebula appears soft and faint, being noticeably affected by the foreground dust absorption, especially in their Western parts (as illustrated in Appendix C).
-Taking this into account, we conclude that the soft X-ray emission can be traced across the full extent of the nebula, except for the “inner cavity” with a radius of pc.
+Since the difference between the lobes’ dimensions amounts to a sizeable fraction of them, one can infer the exponential scale height of the ambient gas density profile of pc, which is consistent with the scale height of the Galactic neutral-gas disk.
+The X-ray emission from the northern and southern boundaries of the nebula appears soft and faint and is noticeably affected by the foreground dust absorption, especially in their western parts (as illustrated in Appendix C).
+Taking this into account, we conclude that the soft X-ray emission can be traced across the full extent of the nebula, except for the “inner cavity,” with a radius of pc.
 The outer boundary of the emission of the northern and southern arcs (outlined as dashed annuli including Region 4 in Fig.
-5) is located at from the center, coinciding with the boundary of the W50’s radio emission.
-This morphology suggests that the soft X-ray emitting gas is confined in a thick () close-to-spherical shell, driving a strong shock in the surrounding ISM.
-The Eastern and the Western lobes are roughly confined inside the projection of the cone with the half-opening angle , while the EXJs encompass only the cone with a twice smaller opening angle,.
-From the eROSITA spectrally decomposed RGB images, a gentle spectral evolution of the EXJs emission with distance is seen clearly.
-EXJ emission is significantly harder at its sharp inner boundary than further away from SS 433, where EXJs blend with softer thermal emission that fills the rest of the lobe, in agreement with earlier findings with XMM-Newton and NuSTAR (Brinkmann et al., 2007; Safi-Harb et al., 2022).
+5) is located at from the center, coinciding with the boundary of W50’s radio emission.
+This morphology suggests that the soft X-ray-emitting gas is confined in a thick () close-to-spherical shell, driving a strong shock in the surrounding ISM.
+The eastern and western lobes are roughly confined inside the projection of the cone with the half-opening angle of , while the EXJs only encompass the cone with an opening angle that is two times smaller, i.e., (see the characteristic angles indicated in Fig.
+5).
+The opening angle of the baryonic jets launched by the central source is much smaller: (e.g., Marshall et al., 2002).
+Therefore, if the jet direction precesses slowly and the ejected matter moves ballistically, it would only fill thin “walls” of a hollow cone, making a tightly wound spiral with cm pc separation ( at kpc).
+Due to a light-travel-time delay between the approaching and receding parts of this spiral, however, the apparent spacing between individual twists should be completely smeared out at distances beyond 1 pc of the central source (e.g., Khabibullin and Sazonov, 2017).
+From the eROSITA spectrally decomposed pseudo-color (red-green-blue) images, a gentle spectral evolution of the EXJs’ emission with distance is seen clearly.
+EXJ emission is significantly harder at its sharp inner boundary than further away from SS 433, where EXJs blend with softer thermal emission that fills the rest of the lobe; this is in agreement with earlier findings of XMM-Newton and NuSTAR (Brinkmann et al., 2007; Safi-Harb et al., 2022).
 5 X-ray spectra
@@
 Since this source is extremely bright, the observations were intentionally performed during the two consecutive orbital eclipses of the compact object (and the jets) by the companion star, so its soft X-ray flux was dimmer by a factor of several compared to the typical levels (Kawai et al., 1989; Lopez et al., 2006; Marshall et al., 2013).
-The combined X-ray spectrum of SS 433 extracted from arcmin aperture (with background estimated from to 5 arcmin annuli, being however completely negligible compared the source) using the data of all pointings in which it falls within Field-of-View is shown in Fig.
+The combined X-ray spectrum of SS 433 extracted from an arcmin aperture (with background estimated from to 5 arcmin annuli; this is, however, completely negligible compared the source) using the data of all pointings in which it falls within the field of view is shown in Fig.
 6.
-As expected, it is well described by the model of the optically thin multitemperature emission from collimated flows that adiabatically expand and cool at distances from a few to a few cm from the central object (Brinkmann et al., 1988, 1991, 2005; Kotani et al., 1996; Marshall et al., 2002; Medvedev & Fabrika, 2010; Khabibullin et al., 2016; Medvedev et al., 2019).
-We discuss it more in Appendix A, while the main focus of this study is instead on the much fainter diffuse emission from the surrounding nebula.
+As expected, it is well described by the model of the optically thin multi-temperature emission from collimated flows that adiabatically expand and cool at distances from a few to a few cm from the central object (Brinkmann et al., 1988, 1991, 2005; Kotani et al., 1996; Marshall et al., 2002; Medvedev and Fabrika, 2010; Khabibullin et al., 2016; Medvedev et al., 2019).
+We discuss this further in Appendix A; the main focus of this study is instead on the much fainter diffuse emission from the surrounding nebula.
 5.2 Diffuse emission
 The high sensitivity of the presented data allowed us to perform spectroscopy of the diffuse emission in several regions and determine the physical state of the radiating plasma.
-We have selected eight regions, shown as numbered boxes in Figure 5, for detailed spectral analysis of their X-ray emission.
+We selected eight regions, shown as numbered boxes in Figure 5, for a detailed spectral analysis of their X-ray emission.
 For each region, the background (estimated from the adjacent regions) was subtracted to isolate the net signal.
-We present the parameters of the fitted models in Table 1 and discuss the implications of the corresponding results below.
+We present the parameters of the fit models in Table 4 and discuss the implications of the corresponding results below.
 5.2.1 Eastern lobe
-Emission from the brightest filament in the Eastern lobe (Region 1 in Figure 5) features a spectrum characteristic for a thermal (but non-equilibrium) optically-thin plasma with a temperature of , as evidenced by the line emission from the highly ionized oxygen, neon, magnesium, silicon and iron (see the left panel in Figure 7).
-The line emissivity and ratios are consistent with the Solar abundance of heavy elements, but also indicate a non-equilibrium state of the gas ionization.
-Namely, it is qualitatively consistent with the expectation for a plane-parallel shell of the gas that experienced a passage of the shock with velocity (from the Rankine-Hugoniot relations for the temperature and density jump properties) some yr ago.
+Emission from the brightest filament in the eastern lobe (Region 1 in Figure 5) features a spectrum characteristic for a thermal (but nonequilibrium) optically-thin plasma with a temperature of , as evidenced by the line emission from the highly ionized oxygen, neon, magnesium, silicon, and iron (see the left panel in Figure 7).
+The line emissivity and ratios are consistent with the solar abundance of heavy elements, but they also indicate a nonequilibrium state of the gas ionization.
+In particular, the simultaneous presence of spectral lines from ion species correspond to a broad range of gas temperatures (from helium-like oxygen O VII to numerous iron L-shell ions), likely reflecting an ongoing ionization process of initially cold and rapidly heated ISM gas (although a multi-temperature mixture of the equilibrium spectra is also capable of reproducing similar features).
+The obtained best-fit, nonequillibrium spectral model is qualitatively consistent with the expectation for a plane-parallel shell of the gas that experienced a passage of the shock with velocity, (from the Rankine-Hugoniot relations for the temperature and density jump properties) yr ago (taking mean molecular weight ).
 The inferred hydrogen number density, cm-3, is a factor of a few higher than the expected gas density of the interstellar gas pc away from the plane of the Galactic disk.
-The spectrum of the diffuse emission in the Eastern lobe (exemplified by Region 3 in Figure 5) also can be well described by thermal non-equilibrium emission of the gas with somewhat higher temperature keV and lower hydrogen number density cm-3 (see the left panel in Figure 7).
+The spectrum of the diffuse emission in the eastern lobe (exemplified by Region 3 in Figure 5) also can be well described by thermal nonequilibrium emission of the gas with a somewhat higher temperature of keV and lower hydrogen number density of cm-3 (see the left panel in Figure 7).
 5.2.2 Western lobe
 The emission from the western lobe appears to be noticeably harder, partially due to the higher interstellar absorption in that direction.
-However, it also corresponds to the emission of plasma with genuinely higher temperature, keV, and a similar hydrogen number density, cm-3 (and a few times higher for the brightest filament), as illustrated by spectra from Regions 6 and 7 in Figure 5).
-The non-equilibrium ionization model (see the right panel in Figure 7) results in a comparable estimate of yr for the ionization time-scale.
-The total absorption-corrected luminosity of the diffuse soft X-ray emission from the nebula is in 0.5-4 keV, which is by a factor of few tens smaller than the apparent X-ray luminosity of the central object, few ,222Due to the beaming of the radiation along the thick accretion disc axis (expected in the case of highly supercritical accretion), the actual X-ray luminosity of SS 433 might be orders of magnitude larger, namely comparable to the Eddington luminosity and kinetic luminosity of the jets, (e.g.
-Fabrika & Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007; Medvedev & Fabrika, 2010; Khabibullin & Sazonov, 2016; Middleton et al., 2021; Fogantini et al., 2023).
-and much smaller than the supposed kinetic luminosity of the outflows from the supercritical accretion disk, .
-However, the thermal energy content of the X-ray emitting gas in the Eastern lobe (with estimated gas number density 0.1 cm-3, temperature 0.5 keV and volume assuming roughly to cylindrical geometry ) is erg, requiring a supply of at least over years.
+However, it also corresponds to the emission of plasma with genuinely higher temperature, keV, and a similar hydrogen number density, cm-3 (this is a few times higher for the brightest filament), as illustrated by spectra from Regions 6 and 7 in Figure 5).
+The nonequilibrium ionization model (see the right panel in Figure 7) results in a comparable estimate of yr for the ionization timescale.
+The total absorption-corrected luminosity of the diffuse soft X-ray emission from the nebula is in 0.5-4 keV, which is several tens times smaller than the apparent X-ray luminosity of the central object, few ,222Due to the beaming of the radiation along the thick-accretion-disk axis (expected in the case of highly supercritical accretion), the actual X-ray luminosity of SS 433 might be orders of magnitude larger and namely comparable to the Eddington luminosity and kinetic luminosity of the jets: (e.g., Fabrika and Mescheryakov, 2001; Begelman et al., 2006; Poutanen et al., 2007; Medvedev and Fabrika, 2010; Khabibullin and Sazonov, 2016; Middleton et al., 2021; Fogantini et al., 2023).
+and much smaller than the supposed kinetic luminosity of the outflows from the supercritical accretion disk: .
+However, the thermal-energy content of the X-ray-emitting gas in the eastern lobe (with an estimated gas number density of 0.1 cm-3, temperature of 0.5 keV, and volume of , assuming roughly cylindrical geometry) is erg, requiring a supply of at least over years.
 The mass of this gas amounts to , implying that the bulk of it cannot be provided by the outflows from SS 433.
-If the volume filling factor of the emitting medium is, however, low, the estimate for the required mass and energy could be decreased by a factor of , but similarly, the density should be increased, leading to the corresponding decrease in the time since the shock passage.
-That means that the X-ray emitting gas provides an estimate of the average energy input in the lobes at the level of , which is comparable to the kinetic luminosity of the jets and the Eddington luminosity of the system (for the mass of the compact object below 10).
+If the volume filling factor, of the emitting medium is low, however, the estimate for the required mass and energy could be decreased by a factor of , but similarly, the density should be increased, leading to the corresponding decrease in the time since the shock passage.
+That means that the X-ray-emitting gas provides an estimate of the average energy input in the lobes at the level of , which is comparable to the kinetic luminosity of the jets and the Eddington luminosity of the system (for the mass of the compact object below 10).
 A similar estimate can be obtained from the western lobe, implying nearly symmetric energy injection along the nebula’s axis at a very high rate over the last few tens of thousands of years.
-Given that the temperature inferred from the spectral fits with the non-equilbirum emission is driven mostly by characteristics of the ionization state, it is likely a lower limit on the actual temperature of the emitting gas.
-Hence, a significant portion of the gas energy content might be effectively hidden, while the required energy input should be increased correspondingly.
+Given that the temperature inferred from the spectral fits with the nonequilbirum emission is driven mostly by characteristics of the ionization state, it is likely a lower limit on the actual temperature of the emitting gas.
+Hence, a significant portion of the gas energy content might be effectively hidden, while the required energy input should be increased accordingly.
 5.2.3 Northern and southern boundaries
 The emission from the nebula boundaries is exemplified by the northern arc region and the southern extension (Regions 4 and 8 in Figure 5).
-The thermal non-equilibrium model provides the temperature estimate keV, low hydrogen number density, cm-3 for the Northern Arc and even lower for the Southern Extension.
-The ionization timescale is, however, similar here as well, yr.
-This observation contrasts with the brightness of the radio emission, which in the Northern Arc region is one of the brightest across the nebula.
+The thermal-nonequilibrium model provides a temperature estimate of keV and low hydrogen number density of cm-3 for the northern arc, and this is even lower for the southern extension.
+The ionization timescale is, however, similar: yr.
+The derived parameters of the nonequillibrium ionization model for the northern-arc region (listed in Table 4) are in good agreement with the findings of Chi et al.
+(2024), which studied a similar region using the data of dedicated XMM-Newton observations guided by earlier mosaic of the whole W50 nebula by ROSAT PSPC (Brinkmann et al., 1996).
+A great advantage of the SRG/eROSITA data (particularly important for the very faint extended emission from these regions) is significantly more robust control over contributions of the instrumental and astrophysical backgrounds.
+This allows much tighter constraints to be obtained on the key parameters of interest here.
+The observation of rather faint X-ray emission contrasts with the brightness of the radio emission, which in the northern-arc region is one of the brightest across the nebula.
 The high degree of linear polarization of the radio emission might indicate that the preferential orientation of the interstellar magnetic field in this region favors efficient acceleration of relativistic particles producing synchrotron emission.
-We note here that the central part of the W50 X-ray map suffers from higher photoelectric absorption than the Eastern and even Western lobes, as illustrated in Fig.
+We note that the central part of the W50 X-ray map suffers from higher photoelectric absorption than the eastern and even western lobes, as illustrated in Fig.
 9.
@@
 5.2.4 Extended X-ray jets
-In contrast to this “thermal” picture, the X-ray spectra of X-ray emission from the “extended X-ray jets” demonstrate a featureless spectrum that can be described by a power law with the slope in the range from (at their base, e.g., Region 5 in Fig.
-5) to (at the termination points, e.g.
+In contrast to this “thermal” picture, the X-ray spectra of X-ray emission from the EXJs demonstrate a featureless spectrum that can be described by a power law with a slope ranging from (at their base; e.g., Region 5 in Fig.
+5) to (at the termination points; e.g.
 Region 2 in Fig.
 5).
-In agreement with previous spectral studies of this emission (e.g., Brinkmann et al., 2007), the thermal bremsstrahlung model also provides an equally good fit with temperatures from keV (at the termination) to keV (at the base).
-The gas emissivity required to produce the X-ray luminosity at level implies hydrogen number densities cm-3 in these regions (assuming a filling factor of unity).
-Clearly, if the thermal nature of this emission were true, it would have a drastically higher pressure compared to the ambient soft X-ray emitting gas and would expand sideways supersonically, driving a strong shock wave.
+This picture is consistent with the results of focused spectral studies of the western EXJ, including the data above 10 keV from NuSTAR observations (Safi-Harb et al., 2022).
+Also, in agreement with previous spectral studies of this emission (e.g., Brinkmann et al., 2007), the thermal bremsstrahlung model also provides an equally good fit with temperatures from keV (at the termination) to keV (at the base).
+The gas emissivity required to produce the X-ray luminosity at level implies hydrogen number densities of cm-3 in these regions (assuming a filling factor of unity).
+Clearly, if the thermal nature of this emission were true, it would have a drastically higher pressure compared to the ambient, soft X-ray-emitting gas and would expand sideways supersonically, driving a strong shock wave.
 Alternatively, the emission might result from the population of relativistic particles accelerated within EXJs.
@@
 S.
-Collaboration et al., 2024) strongly supports the non-thermal scenario.
+Collaboration et al., 2024) strongly supports the nonthermal scenario.
 The absence of any noticeable radio, infrared, or optical emission from these structures would then require substantial self-absorption of radiation at these wavelengths (which is unlikely given that the gas number density is likely very low inside them) or a rather peculiar energy distribution of the relativistic particles with a large fraction of them having energies in excess of tens of TeV.
-Another possibility could be scattering of very bright and beamed emission of the central source (e.g., Panferov & Kabanov, 2013; Khabibullin & Sazonov, 2016, 2019), but the small optical depth with respect to the Thomson scattering, , and the observed spectral gradient along these structures disfavor this possibility.
+Another possibility could be scattering of very bright and beamed emission of the central source (e.g., Panferov and Kabanov, 2013; Khabibullin and Sazonov, 2016, 2019), but the small optical depth with respect to the Thomson scattering, , and the observed spectral gradient along these structures disfavor this possibility.
 6 Discussion
-The Eastern EXJ was well mapped in X-rays with XMM-Newton, Chandra, and NuSTAR (e.g., Brinkmann et al., 2007; Safi-Harb et al., 2022), revealing its non-thermal nature.
+The eastern EXJ was well mapped in X-rays with XMM-Newton, Chandra, and NuSTAR (e.g., Brinkmann et al., 2007; Safi-Harb et al., 2022), revealing its nonthermal nature.
 Furthermore, recent IXPE observation (Kaaret et al., 2024) revealed polarized X-ray emission from a patch adjacent to the base of the eastern EXJ.
 These data suggest that X-ray photons are produced by a synchrotron mechanism with the magnetic field predominantly aligned with the EXJ axis.
-The eROSITA data show that the bases of both EXJs are located at essentially the same distances from SS 433, have similarly sharp inner boundaries, and non-thermal (featureless) spectra across their entire extents.
+The eROSITA data show that the bases of both EXJs are located at essentially the same distances from SS 433 and have similarly sharp inner boundaries and nonthermal (featureless) spectra across their entire extents.
 These structures have often been considered as a direct consequence of the relativistic jets’ propagation inside the nebula before they deposit the bulk of their kinetic energy in the lobes, energizing them and the whole extent of the nebula (e.g., Panferov, 2014).
-However, the direction of the SS 433 jets is known to precess about the mean axis with the amplitude degrees, which is a factor of 2 larger than the observed opening angle of EXJ.
-Since this extended X-ray emission comes from distances of tens of pc, i.e.
-orders of magnitude further away from the central source than the compact radio and X-ray jets, which are traced up to a few cm, the actual connection between them is not obvious (e.g., Panferov, 2017).
-It might be governed by recollimation of the jets’ flows at some intermediate distance from the source (Eichler, 1983; Peter & Eichler, 1993), a historical change in the jets’ direction (Goodall et al., 2011a; Monceau-Baroux et al., 2015), or a different regime of baryonic jets propagation through the magnetized ISM after their gas “freezes out” after recombining due to rapid adiabatic expansion, cooling losses and lack of ambient interactions (Churazov et al., 2020).
-If we assume that there were no recent drastic changes in the properties of W50/SS 433 (e.g., temporary quenching of the central source), the overall picture of the “large-scale energy flow” in the W50/SS 433, supported by X-ray data, can be schematically decomposed into three major components (illustrated in Fig.
-8):
-“Dark flow”: from a compact source up to the base of EXJs covering the range of distances from to from SS 433.
+However, the direction of the SS 433 jets is known to precess at the mean axis with an amplitude of degrees, which is a factor of two larger than the observed opening angle of an EXJ.
+Since this extended X-ray emission comes from distances of approximately tens of parsec; i.e., orders of magnitude further away from the central source than the compact radio and X-ray jets, which are traced up to a few cm, the actual connection between them is not obvious (e.g., Panferov, 2017).
+It might be governed by recollimation of the jets’ flows at a given intermediate distance from the source (Eichler, 1983; Peter and Eichler, 1993), a historical change in the jets’ direction (Goodall et al., 2011a; Monceau-Baroux et al., 2015), or a different regime of baryonic jet propagation through the magnetized interstellar medium (ISM) after their gas “freezes out” following recombination due to rapid adiabatic expansion, cooling losses, and lack of ambient interactions (Churazov et al., 2020).
+If we assume that there were no recent drastic changes in the properties of W50/SS 433 (e.g., temporary quenching of the central source), the overall picture of the “large-scale energy flow” in W50/SS 433, supported by X-ray data, can be schematically decomposed into the three major components (illustrated in Fig.
+8) listed below.
+Dark flow: this spreads from a compact source up to the base of EXJs covering the range of distances from to from SS 433.
 Over this distance range, there are no visible signatures in the X-ray (or any other) band.
-“Non-thermal flow”: from the base to the end of EXJs that have featureless non-thermal spectra (presumably, synchrotron emission).
-“Thermal flow”: extended diffuse regions between the W50 radio boundaries and EXJs. Its X-ray emission is typical for ISM shock-heated to sub-keV temperatures.
-Analytical and numerical models primarily aimed at reproducing W50’s radio morphology often assume that a quasi-spherical part of the nebula is a remnant of the supernova explosion, while the elongated structures are directly associated with the energy release by the binary system (e.g., Begelman et al., 1980; Zealey et al., 1980; Eichler, 1983; Peter & Eichler, 1993; Velázquez & Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Asahina et al., 2014; Monceau-Baroux et al., 2015; Panferov, 2017; Ohmura et al., 2021).
+Nonthermal flow: this is from the base to the end of EXJs that have featureless nonthermal spectra (presumably synchrotron emission).
+Thermal flow: this represents extended diffuse regions between the W50 radio boundaries and EXJs. Its X-ray emission is typical for an ISM that is shock-heated to sub-keV temperatures.
+Analytical and numerical models primarily aimed at reproducing W50’s radio morphology often assume that a quasi-spherical part of the nebula is a remnant of the supernova explosion, while the elongated structures are directly associated with the energy release by the binary system (e.g., Begelman et al., 1980; Zealey et al., 1980; Eichler, 1983; Peter and Eichler, 1993; Velázquez and Raga, 2000; Zavala et al., 2008; Goodall et al., 2011a; Asahina et al., 2014; Monceau-Baroux et al., 2015; Panferov, 2017; Ohmura et al., 2021).
 In the model considered by Churazov et al.
 (2024), the entire nebula is powered by the central source.
-There, an anisotropic outflow/wind from the Hyper-Eddington accretor is assumed - a combination of a more powerful collimated outflow along the orbital axis and a quasi-isotropic wind in all other directions.
-In this model, the “Dark flow” phase corresponds to the free expansion of the wind, and the onset of the EXJs is associated with the termination shock of the isotropic wind, which initiates shocks in the collimated flow where the pressure rises suddenly.
+There, an anisotropic outflow/wind from the hyper-Eddington accretor is assumed: a combination of a more powerful collimated outflow along the orbital axis and a quasi-isotropic wind in all other directions.
+In this model, the darkflow phase corresponds to the free expansion of the wind, and the onset of the EXJs is associated with the termination shock of the isotropic wind, which initiates shocks in the collimated flow where the pressure rises suddenly.
 Finally, the thermal part is the shock-heated ISM.
 In this model, the characteristic shape of the W50 nebula (a quasi-spherical part and two extensions) reflects the anisotropy of the velocity and kinetic power of the wind, with no need for the impact of the narrow baryonic jets at all.
-In all these models, the outer shock, which delineates the radio boundary of W50, is not dissimilar to the shocks seen around middle-aged supernova remnants or stellar wind-blown bubbles (e.g., Konigl, 1983; Dubner et al., 1998).
+In all these models, the outer shock, which delineates the radio boundary of W50, is not dissimilar to the shocks seen around middle-aged SNRs or stellar wind-blown bubbles (e.g., Konigl, 1983; Dubner et al., 1998).
 What makes W50 “special” is the presence of EXJs, which are radio-faint but bright in X-rays and VHE gamma-rays.
@@
 Bykov et al., 2025) .
-Remarkable symmetry of the EXJs, very clearly demonstrated by the eROSITA image, points towards irrelevance of the ambient ISM properties to the energy and momentum propagation, as well as the particle acceleration regime, meaning that a cocoon or a wind-blown cavity likely encompasses these structures in the current epoch.
-In this regard, the presence of H emitting filaments in several places across the nebula might help probe the energetic content of the nebula (e.g., Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017).
+The remarkable symmetry of the EXJs – which is very clearly demonstrated by the eROSITA image– points toward the irrelevance of the ambient ISM properties to the energy and momentum propagation, as well as the particle acceleration regime, meaning that a cocoon or a wind-blown cavity likely encompasses these structures in the current epoch.
+In this regard, the presence of H-emitting filaments in several places across the nebula might help probe the energetic content of the nebula (e.g., Boumis et al., 2007; Abolmasov et al., 2010; Farnes et al., 2017).
 Interestingly, the location of the brightest H filaments does not correspond to any prominent features in thermal X-ray emission.
@@
 Given the presence of bright silicon lines in the spectra extracted from these regions, an observation of them with the XRISM observatory (Tashiro et al., 2025) might provide estimates of the directed and turbulent velocities.
-A large grasp soft X-ray microcalorimeter mission, like LEM (Kraft et al., 2022) or HUBS (Bregman et al., 2023) would also be perfectly suited for in-depth spectral exploration of this complex system with multiple interacting components.
+A large grasp soft X-ray microcalorimeter mission such as LEM (Kraft et al., 2022) or HUBS (Bregman et al., 2023) would also be perfectly suited for the in-depth spectral exploration of this complex system with multiple interacting components.
 7 Conclusions
-SRG/eROSITA observations of W50/SS 433 provide, for the first time, the entire X-ray map of the nebula with high spatial and spectral resolution.
-The data support a physical picture in which the anisotropic energy flow from SS 433 evolves passing through three distinct stages:
-an invisible “dark” flow between 0.1 and 25 pc (=anisotropic wind from the binary system)
-a “non-thermal” flow over another (=EXJs)
-a thermal flow (= shock-heated ISM) that envelopes EXJs.
+The SRG/eROSITA observations of W50/SS 433 provide, for the first time, the entire X-ray map of the nebula with high spatial and spectral resolution.
+The data support a physical picture in which the anisotropic energy flow from SS 433 evolves passing through the following three distinct stages:
+an invisible dark flow between 0.1 and 25 pc (=anisotropic wind from the binary system)
+a nonthermal flow over another (=EXJs)
+a thermal flow (= shock-heated ISM) that envelopes EXJs
 The appearance of the nebula is further affected by the large ambient density gradients and a heavy foreground photoelectric absorption that projects almost exactly on the central quasi-spherical part of the W50 nebula.
 The thermal part of the W50 X-ray emission can be reasonably well described by the shock-heated plasma that has not yet reached temperature and ionization equilibrium.
-Such emission is typical for middle-age or old SNRs.
+Such emission is typical for middle-aged or old SNRs.
 The outer radio boundary of the nebula is also reminiscent of the SNR shocks.
-On the contrary, the “Extended X-ray Jets” are the most remarkable features of this system on scales of tens of pc.
-Their sharp inner edges plausibly correspond to extreme shocks that accelerate particles and power the X-ray (synchrotron) emission, and, at 10 orders of magnitude higher energies, the TeV emission.
-W50/SS 433 system illustrates clearly the important role the Hyper-Eddington accretors might play for the energetics of the ISM in galaxies at different redshifts and the production of ultra-high-energy particles.
+On the contrary, the EXJs are the most remarkable features of this system on scales of tens of parsec .
+Their sharp inner edges plausibly correspond to extreme shocks that accelerate particles and power the X-ray (synchrotron) emission, and, at energy levels that are ten orders of magnitude higher, the TeV emission.
+The W50/SS 433 system clearly illustrates the important role the hyper-Eddington accretors might play in the energetics of the ISM in galaxies at different redshifts and the production of ultra-high-energy particles.
+Data availability
+X-ray data analyzed in this article were used with the permission of the Russian SRG/eROSITA consortium.
+The reduced broad-band image (Fig.
+2) is available at the CDS via anonymous ftp to cdsarc.cds.unistra.fr (130.79.128.5).
 Acknowledgements.
@@
 The eROSITA data were processed using the eSASS/NRTA software system developed by the German eROSITA consortium and analyzed using proprietary data reduction software developed by the Russian eROSITA Consortium.
-IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.References
-- Abell & Margon (1979) Abell, G. O. & Margon, B. 1979, Nature, 279, 701
-- Abeysekara et al. (2018) Abeysekara, A. U., Albert, A., Alfaro, R., et al. 2018, Nature, 562, 82
-- Abolmasov et al. (2010) Abolmasov, P., Maryeva, O., & Burenkov, A. N. 2010, AN, 331, 412
-- Alfaro et al. (2024) Alfaro, R., Alvarez, C., Arteaga-Velázquez, J. C., et al. 2024, ApJ, 976, 30
-- Arnason et al. (2021) Arnason, R. M., Papei, H., Barmby, P., Bahramian, A., & Gorski, M. D. 2021, MNRAS, 502, 5455
-- Arnaud (1996) Arnaud, K. A. 1996, in Astronomical Society of the Pacific Conference Series, Vol. 101, Astronomical Data Analysis Software and Systems V, ed. G. H. Jacoby & J. Barnes, 17
-- Asahina et al. (2014) Asahina, Y., Ogawa, T., Kawashima, T., et al. 2014, ApJ, 789, 79
-- Atapin & Fabrika (2016) Atapin, K. E. & Fabrika, S. N. 2016, AstL, 42, 517
-- Begelman et al. (2006) Begelman, M. C., King, A. R., & Pringle, J. E. 2006, MNRAS, 370, 399
-- Begelman et al. (1980) Begelman, M. C., Sarazin, C. L., Hatchett, S. P., McKee, C. F., & Arons, J. 1980, ApJ, 238, 722
-- Blundell & Bowler (2004) Blundell, K. M. & Bowler, M. G. 2004, ApJL, 616, L159
-- Blundell et al. (2011) Blundell, K. M., Schmidtobreick, L., & Trushkin, S. 2011, MNRAS, 417, 2401
-- Borkowski et al. (2001) Borkowski, K. J., Lyerly, W. J., & Reynolds, S. P. 2001, ApJ, 548, 820
-- Boumis et al. (2007) Boumis, P., Meaburn, J., Alikakos, J., et al. 2007, MNRAS, 381, 308
-- Bregman et al. (2023) Bregman, J., Cen, R., Chen, Y., et al. 2023, SCPMA, 66, 299513
-- Brinkmann et al. (1996) Brinkmann, W., Aschenbach, B., & Kawai, N. 1996, A&A, 312, 306
-- Brinkmann et al. (1988) Brinkmann, W., Fink, H. H., Massaglia, S., Bodo, G., & Ferrari, A. 1988, A&A, 196, 313
-- Brinkmann & Kawai (2000) Brinkmann, W. & Kawai, N. 2000, A&A, 363, 640
-- Brinkmann et al. (1991) Brinkmann, W., Kawai, N., Matsuoka, M., & Fink, H. H. 1991, A&A, 241, 112
-- Brinkmann et al. (2005) Brinkmann, W., Kotani, T., & Kawai, N. 2005, A&A, 431, 575
-- Brinkmann et al. (2007) Brinkmann, W., Pratt, G. W., Rohr, S., Kawai, N., & Burwitz, V. 2007, A&A, 463, 611
-- Broderick et al. (2018) Broderick, J. W., Fender, R. P., Miller-Jones, J. C. A., et al. 2018, MNRAS, 475, 5360
-- Bykov et al. (2025) Bykov, A. M., Osipov, S. M., Romansky, V. I., et al. 2025, PhRvD, 112, 063017
-- Cherepashchuk et al. (2023) Cherepashchuk, A., Belinski, A., Dodin, A., & Postnov, K. 2023, NewA, 103, 102060
-- Cherepashchuk (1981) Cherepashchuk, A. M. 1981, MNRAS, 194, 761
-- Cherepashchuk et al. (2025) Cherepashchuk, A. M., Dodin, A. V., & Postnov, K. A. 2025, arXiv, arXiv:2506.01106
-- Cherepashchuk et al. (2018) Cherepashchuk, A. M., Esipov, V. F., Dodin, A. V., Davydov, V. V., & Belinskii, A. A. 2018, ARep, 62, 747
-- Chi et al. (2024) Chi, Y.-H., Huang, J., Zhou, P., et al. 2024, ApJL, 975, L28
-- Churazov et al. (2023) Churazov, E., Khabibullin, I., Bykov, A. M., Lyskova, N., & Sunyaev, R. 2023, A&A, 670, A156
-- Churazov et al. (2021a) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Bykov, A. M. 2021a, A&A, 651, A41
-- Churazov et al. (2025) Churazov, E., Khabibullin, I., Lyskova, N., Sunyaev, R., & Dolag, K. 2025, arXiv, arXiv:2507.19987
-- Churazov et al. (2020) Churazov, E., Khabibullin, I., & Sunyaev, R. 2020, MNRAS, 495, L51
-- Churazov et al. (2024) Churazov, E. M., Khabibullin, I. I., & Bykov, A. M. 2024, A&A, 688, A4
-- Churazov et al. (2021b) Churazov, E. M., Khabibullin, I. I., Bykov, A. M., et al. 2021b, MNRAS, 507, 971
-- Dubner et al. (1998) Dubner, G. M., Holdaway, M., Goss, W. M., & Mirabel, I. F. 1998, AJ, 116, 1842
-- Eichler (1983) Eichler, D. 1983, ApJ, 272, 48
-- Eikenberry et al. (2001) Eikenberry, S. S., Cameron, P. B., Fierce, B. W., et al. 2001, ApJ, 561, 1027
-- Elston & Baum (1987) Elston, R. & Baum, S. 1987, AJ, 94, 1633
-- Fabian & Rees (1979) Fabian, A. C. & Rees, M. J. 1979, MNRAS, 187, 13P
-- Fabrika (2004) Fabrika, S. 2004, ASPRv, 12, 1
-- Fabrika & Mescheryakov (2001) Fabrika, S. & Mescheryakov, A. 2001, in IAU Symposium, Vol. 205, Galaxies and their Constituents at the Highest Angular Resolutions, ed. R. T. Schilizzi, 268
-- Farnes et al. (2017) Farnes, J. S., Gaensler, B. M., Purcell, C., et al. 2017, MNRAS, 467, 4777
-- Filippova et al. (2006) Filippova, E., Revnivtsev, M., Fabrika, S., Postnov, K., & Seifina, E. 2006, A&A, 460, 125
-- Fogantini et al. (2023) Fogantini, F. A., García, F., Combi, J. A., et al. 2023, A&A, 669, A149
-- Forman et al. (1978) Forman, W., Jones, C., Cominsky, L., et al. 1978, ApJS, 38, 357
-- Fragile et al. (2025) Fragile, P. C., Middleton, M. J., Bollimpalli, D. A., & Smith, Z. 2025, MNRAS, 540, 2820
-- Fuchs et al. (2006) Fuchs, Y., Koch Miramond, L., & Ábrahám, P. 2006, A&A, 445, 1041
-- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
-- Geldzahler et al. (1980) Geldzahler, B. J., Pauls, T., & Salter, C. J. 1980, A&A, 84, 237
-- Goodall et al. (2011a) Goodall, P. T., Alouani-Bibi, F., & Blundell, K. M. 2011a, MNRAS, 414, 2838
-- Goodall et al. (2011b) Goodall, P. T., Blundell, K. M., & Bell Burnell, S. J. 2011b, MNRAS, 414, 2828
-- Goranskij (2011) Goranskij, V. 2011, PZ, 31, 5
-- Green (2025) Green, D. A. 2025, JApA, 46, 14
-- Green et al. (2019) Green, G. M., Schlafly, E., Zucker, C., Speagle, J. S., & Finkbeiner, D. 2019, ApJ, 887, 93
-- Gregory et al. (1996) Gregory, P. C., Scott, W. K., Douglas, K., & Condon, J. J. 1996, ApJS, 103, 427
-- Grindlay et al. (1984) Grindlay, J. E., Band, D., Seward, F., et al. 1984, ApJ, 277, 286
-- H. E. S. S. Collaboration et al. (2024) H. E. S. S. Collaboration, Aharonian, F., Ait Benkhali, F., et al. 2024, Sci, 383, 402
-- Heinz & Sunyaev (2002) Heinz, S. & Sunyaev, R. 2002, A&A, 390, 751
-- Hjellming & Johnston (1981a) Hjellming, R. M. & Johnston, K. J. 1981a, ApJL, 246, L141
-- Hjellming & Johnston (1981b) Hjellming, R. M. & Johnston, K. J. 1981b, Nature, 290, 100
-- Holden & Caswell (1969) Holden, D. J. & Caswell, J. L. 1969, MNRAS, 143, 407
-- Jiang et al. (2019) Jiang, Y.-F., Stone, J. M., & Davis, S. W. 2019, ApJ, 880, 67
-- Kaaret et al. (2017) Kaaret, P., Feng, H., & Roberts, T. P. 2017, ARA&A, 55, 303
-- Kaaret et al. (2024) Kaaret, P., Ferrazzoli, R., Silvestri, S., et al. 2024, ApJL, 961, L12
-- Katz (1980) Katz, J. I. 1980, ApJL, 236, L127
-- Katz et al. (1982) Katz, J. I., Anderson, S. F., Margon, B., & Grandi, S. A. 1982, ApJ, 260, 780
-- Kawai et al. (1989) Kawai, N., Matsuoka, M., Pan, H.-C., & Stewart, G. C. 1989, PASJ, 41, 491
-- Khabibullin et al. (2022) Khabibullin, I., Churazov, E., & Sunyaev, R. 2022, MNRAS, 509, 6068
-- Khabibullin et al. (2016) Khabibullin, I., Medvedev, P., & Sazonov, S. 2016, MNRAS, 455, 1414
-- Khabibullin & Sazonov (2016) Khabibullin, I. & Sazonov, S. 2016, MNRAS, 457, 3963
-- Khabibullin et al. (2023) Khabibullin, I. I., Churazov, E. M., Bykov, A. M., Chugai, N. N., & Sunyaev, R. A. 2023, MNRAS, 521, 5536
-- Khabibullin et al. (2024) Khabibullin, I. I., Churazov, E. M., Chugai, N. N., et al. 2024, A&A, 689, A278
-- Khabibullin & Sazonov (2012) Khabibullin, I. I. & Sazonov, S. Y. 2012, AstL, 38, 443
-- Khabibullin & Sazonov (2017) Khabibullin, I. I. & Sazonov, S. Y. 2017, AstL, 43, 388
-- Khabibullin & Sazonov (2019) Khabibullin, I. I. & Sazonov, S. Y. 2019, AstL, 45, 282
-- Kimura et al. (2020) Kimura, S. S., Murase, K., & Mészáros, P. 2020, ApJ, 904, 188
-- Kirshner & Chevalier (1980) Kirshner, R. P. & Chevalier, R. A. 1980, ApJL, 242, L77
-- Konigl (1983) Konigl, A. 1983, MNRAS, 205, 471
-- Kotani et al. (1996) Kotani, T., Kawai, N., Matsuoka, M., & Brinkmann, W. 1996, PASJ, 48, 619
-- Kraft et al. (2022) Kraft, R., Markevitch, M., Kilbourne, C., et al. 2022, arXiv, arXiv:2211.09827
-- Kubota et al. (2010) Kubota, K., Ueda, Y., Kawai, N., et al. 2010, PASJ, 62, 323
-- Leibowitz (1984) Leibowitz, E. M. 1984, MNRAS, 210, 279
-- LHAASO Collaboration (2024) LHAASO Collaboration. 2024, arXiv, arXiv:2410.08988
-- Lopez et al. (2006) Lopez, L. A., Marshall, H. L., Canizares, C. R., Schulz, N. S., & Kane, J. F. 2006, ApJ, 650, 338
-- Margon (1984) Margon, B. 1984, ARA&A, 22, 507
-- Margon et al. (1979) Margon, B., Ford, H. C., Katz, J. I., et al. 1979, ApJL, 230, L41
-- Marshall et al. (2013) Marshall, H. L., Canizares, C. R., Hillwig, T., et al. 2013, ApJ, 775, 75
-- Marshall et al. (2002) Marshall, H. L., Canizares, C. R., & Schulz, N. S. 2002, ApJ, 564, 941
-- Medvedev & Fabrika (2010) Medvedev, A. & Fabrika, S. 2010, MNRAS, 402, 479
-- Medvedev et al. (2019) Medvedev, P. S., Khabibullin, I. I., & Sazonov, S. Y. 2019, AstL, 45, 299
-- Medvedev et al. (2018) Medvedev, P. S., Khabibullin, I. I., Sazonov, S. Y., Churazov, E. M., & Tsygankov, S. S. 2018, AstL, 44, 390
-- Medvedev et al. (2022) Medvedev, P. S., Khabibullin, I. I., Semena, A. N., et al. 2022, AstL, 48, 389
-- Middleton et al. (2021) Middleton, M. J., Walton, D. J., Alston, W., et al. 2021, MNRAS, 506, 1045
-- Migliari et al. (2005) Migliari, S., Fender, R. P., Blundell, K. M., Méndez, M., & van der Klis, M. 2005, MNRAS, 358, 860
-- Miller-Jones et al. (2008) Miller-Jones, J. C. A., Migliari, S., Fender, R. P., et al. 2008, ApJ, 682, 1141
-- Monceau-Baroux et al. (2015) Monceau-Baroux, R., Porth, O., Meliani, Z., & Keppens, R. 2015, A&A, 574, A143
-- Namiki et al. (2000) Namiki, M., Kawai, N., Kotani, T., Mamauchi, S., & Brinkmann, W. 2000, AdSpR, 25, 709
-- Ohmura et al. (2021) Ohmura, T., Ono, K., Sakemi, H., et al. 2021, ApJ, 910, 149
-- Ohsuga & Mineshige (2011) Ohsuga, K. & Mineshige, S. 2011, ApJ, 736, 2
-- Panferov (2014) Panferov, A. 2014, A&A, 562, A130
-- Panferov (2017) Panferov, A. A. 2017, A&A, 599, A77
-- Panferov & Kabanov (2013) Panferov, A. A. & Kabanov, A. A. 2013, arXiv, arXiv:1306.4486
-- Peretti et al. (2025) Peretti, E., Petropoulou, M., Vasilopoulos, G., & Gabici, S. 2025, A&A, 698, A188
-- Peter & Eichler (1993) Peter, W. & Eichler, D. 1993, ApJ, 417, 170
-- Poutanen et al. (2007) Poutanen, J., Lipunova, G., Fabrika, S., Butkevich, A. G., & Abolmasov, P. 2007, MNRAS, 377, 1187
-- Predehl et al. (2021) Predehl, P., Andritschke, R., Arefiev, V., et al. 2021, A&A, 647, A1
-- Predehl et al. (2020) Predehl, P., Sunyaev, R. A., Becker, W., et al. 2020, Nature, 588, 227
-- Rosado et al. (2021) Rosado, M., Sánchez-Cruces, M., Ambrocio-Cruz, P., & Trejo, D. 2021, MNRAS, 506, 4263
-- Safi-Harb et al. (2022) Safi-Harb, S., Mac Intyre, B., Zhang, S., et al. 2022, ApJ, 935, 163
-- Safi-Harb & Ögelman (1997) Safi-Harb, S. & Ögelman, H. 1997, ApJ, 483, 868
-- Safi-Harb & Petre (1999) Safi-Harb, S. & Petre, R. 1999, ApJ, 512, 784
-- Sakai et al. (2025) Sakai, Y., Yamada, S., Sakemi, H., et al. 2025, arXiv, arXiv:2507.19042
-- Sakemi et al. (2021) Sakemi, H., Omae, R., Ohmura, T., & Machida, M. 2021, PASJ, 73, 530
-- Sarazin et al. (1980) Sarazin, C. L., Begelman, M. C., & Hatchett, S. P. 1980, ApJL, 238, L129
-- Seward et al. (1980) Seward, F., Grindlay, J., Seaquist, E., & Gilmore, W. 1980, Nature, 287, 806
-- Seward et al. (1976) Seward, F. D., Page, C. G., Turner, M. J. L., & Pounds, K. A. 1976, MNRAS, 175, 39P
-- Shakura & Sunyaev (1973) Shakura, N. I. & Sunyaev, R. A. 1973, A&A, 24, 337
-- Sharpless (1959) Sharpless, S. 1959, ApJS, 4, 257
-- Shklovskii (1981) Shklovskii, I. S. 1981, SvA, 25, 315
-- Shuder et al. (1980) Shuder, J. M., Hatfield, B. F., & Cohen, R. D. 1980, PASP, 92, 259
-- Spencer (1979) Spencer, R. E. 1979, Nature, 282, 483
-- Stephenson & Sanduleak (1977) Stephenson, C. B. & Sanduleak, N. 1977, ApJS, 33, 459
-- Stewart et al. (1987) Stewart, G. C., Watson, M. G., Matsuoka, M., et al. 1987, MNRAS, 228, 293
-- Sunyaev et al. (2021) Sunyaev, R., Arefiev, V., Babyshkin, V., et al. 2021, A&A, 656, A132
-- Tashiro et al. (2025) Tashiro, M., Kelley, R., Watanabe, S., et al. 2025, PASJ
-- Tsuji et al. (2025) Tsuji, N., Inoue, Y., Khangulyan, D., et al. 2025, arXiv, arXiv:2510.06431
-- van den Heuvel (1981) van den Heuvel, E. P. J. 1981, VA, 25, 95
-- van den Heuvel et al. (1980) van den Heuvel, E. P. J., Ostriker, J. P., & Petterson, J. A. 1980, A&A, 81, L7
-- van den Heuvel et al. (2017) van den Heuvel, E. P. J., Portegies Zwart, S. F., & de Mink, S. E. 2017, MNRAS, 471, 4256
-- Velázquez & Raga (2000) Velázquez, P. F. & Raga, A. C. 2000, A&A, 362, 780
-- Wang et al. (2025) Wang, J., Reville, B., & Aharonian, F. A. 2025, ApJL, 989, L25
-- Watson et al. (1986) Watson, M. G., Stewart, G. C., Brinkmann, W., & King, A. R. 1986, MNRAS, 222, 261
-- Watson et al. (1983) Watson, M. G., Willingale, R., Grindlay, J. E., & Seward, F. D. 1983, ApJ, 273, 688
-- Westerhout (1958) Westerhout, G. 1958, BAN, 14, 215
-- Whitmire & Matese (1980) Whitmire, D. P. & Matese, J. J. 1980, MNRAS, 193, 707
-- Wilms et al. (2000) Wilms, J., Allen, A., & McCray, R. 2000, ApJ, 542, 914
-- Yamauchi et al. (1994) Yamauchi, S., Kawai, N., & Aoki, T. 1994, PASJ, 46, L109
-- Yoshioka et al. (2022) Yoshioka, S., Mineshige, S., Ohsuga, K., Kawashima, T., & Kitaki, T. 2022, PASJ, 74, 1378
-- Zavala et al. (2008) Zavala, J., Velázquez, P. F., Cerqueira, A. H., & Dubner, G. M. 2008, MNRAS, 387, 839
-- Zealey et al. (1980) Zealey, W. J., Dopita, M. A., & Malin, D. F. 1980, MNRAS, 192, 731
+IK acknowledges support by the COMPLEX project from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program grant agreement ERC-2019-AdG 882679.
+Some of the maps were built using cubehelix colorscheme developed by David Green (Green, 2011).References
+- A kinematic model for SS433. Nature 279 (5715), pp. 701–703. External Links: Document, ADS entry Cited by: §1.
+- Very-high-energy particle acceleration powered by the jets of the microquasar SS 433. Nature 562 (7725), pp. 82–85. External Links: Document, 1810.01892, ADS entry Cited by: §1, §5.2.4.
+- The Eastern filament of W50. AN 331 (4), pp. 412. External Links: Document, 1003.4706, ADS entry Cited by: §1, §6.
+- Spectral Study of Very-high-energy Gamma Rays from SS 433 with HAWC. ApJ 976 (1), pp. 30. External Links: Document, 2410.21796, ADS entry Cited by: §1, §5.2.4.
+- Distances to Galactic X-ray binaries with Gaia DR2. MNRAS 502 (4), pp. 5455–5470. External Links: Document, 2102.02615, ADS entry Cited by: footnote 1.
+- XSPEC: The First Ten Years.
+In Astronomical Data Analysis Software and Systems V, G.
+H.
+Jacoby and J.
+Barnes (Eds.), Astronomical Society of the Pacific Conference Series, Vol.
+101, pp.
+17.
+External Links: ADS entry Cited by: Appendix B.
+- Magnetohydrodynamic Simulations of a Jet Drilling an H I Cloud: Shock Induced Formation of Molecular Clouds and Jet Breakup.
+ApJ 789 (1), pp.
+79.
+External Links: Document, 1407.2381, ADS entry Cited by: §6.
+- X-ray variability of SS 433: Evidence for supercritical accretion. AstL 42 (8), pp. 517–530. External Links: Document, 1607.04185, ADS entry Cited by: §1.
+- The nature of SS433 and the ultraluminous X-ray sources. MNRAS 370 (1), pp. 399–404. External Links: Document, astro-ph/0604497, ADS entry Cited by: §1, footnote 2.
+- Beam models for SS 433.. ApJ 238, pp. 722–730. External Links: Document, ADS entry Cited by: §1, §6.
+- Symmetry in the Changing Jets of SS 433 and Its True Distance from Us. ApJL 616 (2), pp. L159–L162. External Links: Document, astro-ph/0410456, ADS entry Cited by: §1.
+- SS433’s accretion disc, wind and jets: before, during and after a major flare. MNRAS 417 (4), pp. 2401–2410. External Links: Document, 1104.2917, ADS entry Cited by: §1.
+- Supernova Remnants in the Sedov Expansion Phase: Thermal X-Ray Emission. ApJ 548 (2), pp. 820–835. External Links: Document, astro-ph/0008066, ADS entry Cited by: Appendix B.
+- Deep optical observations of the interaction of the SS 433 microquasar jet with the W50 radio continuum shell.
+MNRAS 381 (1), pp.
+308–318.
+External Links: Document, 0707.4243, ADS entry Cited by: §1, §3, §6.
+- Scientific objectives of the Hot Universe Baryon Surveyor (HUBS) mission. SCPMA 66 (9), pp. 299513. External Links: Document, 2307.05672, ADS entry Cited by: §6.
+- ROSAT observations of the W 50/SS 433 system.. A&A 312, pp. 306–316. External Links: ADS entry Cited by: §1, §1, §2, §2, §3, §5.2.3.
+- The thermal evolution of the jets in SS 433.. A&A 196, pp. 313–326. External Links: ADS entry Cited by: Appendix A, §1, §5.1.
+- The X-ray emission of SS 433.. A&A 241, pp. 112. External Links: ADS entry Cited by: §5.1.
+- The jets of SS 433: second order effects. A&A 363, pp. 640–646. External Links: ADS entry Cited by: Appendix A.
+- XMM-Newton observations of SS 433 I. EPIC spectral analysis. A&A 431, pp. 575–586. External Links: Document, ADS entry Cited by: §1, §1, §2, Figure 6, §5.1.
+- XMM-Newton observations of the eastern jet of SS 433. A&A 463 (2), pp. 611–619. External Links: Document, astro-ph/0610781, ADS entry Cited by: §1, §3, §4, §5.2.4, §6.
+- LOFAR 150-MHz observations of SS 433 and W 50. MNRAS 475 (4), pp. 5360–5377. External Links: Document, 1802.03406, ADS entry Cited by: §1.
+- PeV particle acceleration and nonthermal emission in the minimalist model of the extended jets in W50/SS433.
+PhRvD 112 (6), pp.
+063017.
+External Links: Document, 2509.09883, ADS entry Cited by: §1, §6.
+- Spectroscopic Monitoring of SS 433. Stability of Parameters of the Kinematic Model over 40 Years. ARep 62 (11), pp. 747–763. External Links: Document, ADS entry Cited by: §1.
+- SS 433 as an Eclipsing Binary.. MNRAS 194, pp. 761–769. External Links: Document, ADS entry Cited by: §1.
+- Evolutionary increase of the orbital separation and change of the Roche lobe size in SS433. NewA 103, pp. 102060. External Links: Document, 2305.07093, ADS entry Cited by: §1.
+- Unique microquasar SS433: new results, new issues. PhyU 68 (10), pp. 1042–1060. External Links: Document, 2506.01106, ADS entry Cited by: §1.
+- An X-Ray Shell Reveals the Supernova Explosion for Galactic Microquasar SS 433. ApJL 975 (2), pp. L28. External Links: Document, 2410.06510, ADS entry Cited by: §1, §1, §2, §5.2.3.
+- Tempestuous life beyond R500: X-ray view on the Coma cluster with SRG/eROSITA. II. Shock and relic. A&A 670, pp. A156. External Links: Document, 2205.07511, ADS entry Cited by: §2.
+- Tempestuous life beyond R500: X-ray view on the Coma cluster with SRG/eROSITA.
+I.
+X-ray morphology, recent merger, and radio halo connection.
+A&A 651, pp.
+A41.
+External Links: Document, 2012.11627, ADS entry Cited by: §2.
+- Does Nature use neutral beams for interstellar plasma heating around compact objects?. MNRAS 495 (1), pp. L51–L55. External Links: Document, 2002.02027, ADS entry Cited by: §1, §6.
+- SRG/eROSITA discovery of a large circular SNR candidate G116.6-26.1: SN Ia explosion probing the gas of the Milky Way halo?.
+MNRAS 507 (1), pp.
+971–982.
+External Links: Document, 2106.09454, ADS entry Cited by: §2.
+- Minimalist model of the W50/SS433 extended X-ray jet: Anisotropic wind with recollimation shocks. A&A 688, pp. A4. External Links: Document, 2401.14770, ADS entry Cited by: §1, §2, §3, Figure 8, §6.
+- Perseus cluster in its X-ray entirety with SRG/eROSITA et al. I. Merger and Radio-Uroboroses. arXiv, pp. arXiv:2507.19987. External Links: Document, 2507.19987, ADS entry Cited by: §2.
+- A High-Resolution Radio Study of the W50-SS 433 System and the Surrounding Medium. AJ 116 (4), pp. 1842–1855. External Links: Document, ADS entry Cited by: §1, Figure 8, §6.
+- Focusing of high-mach number jets by an ambient medium.. ApJ 272, pp. 48–53. External Links: Document, ADS entry Cited by: §6, §6.
+- Twenty Years of Timing SS 433. ApJ 561 (2), pp. 1027–1033. External Links: Document, astro-ph/0107296, ADS entry Cited by: §2.
+- High-Resolution Radio Observations of W50, the Remnant Associated with SS 433. AJ 94, pp. 1633. External Links: Document, ADS entry Cited by: §1.
+- SS 433: a double jet in action?. MNRAS 187, pp. 13P–16. External Links: Document, ADS entry Cited by: §1.
+- Face-on SS 433 stars as a possible new type of extragalactic X-ray sources.
+In Galaxies and their Constituents at the Highest Angular Resolutions, R.
+T.
+Schilizzi (Ed.), IAU Symposium, Vol.
+205, pp.
+268.
+External Links: Document, astro-ph/0103070, ADS entry Cited by: §1, footnote 2.
+- The jets and supercritical accretion disk in SS433. ASPRv 12, pp. 1–152. External Links: Document, astro-ph/0603390, ADS entry Cited by: §1.
+- Interacting large-scale magnetic fields and ionized gas in the W50/SS433 system. MNRAS 467 (4), pp. 4777–4801. External Links: Document, 1604.06552, ADS entry Cited by: §1, §3, §6.
+- Diagnostics of SS433 with the RXTE. A&A 460 (1), pp. 125–131. External Links: Document, astro-ph/0609367, ADS entry Cited by: §1.
+- A NuSTAR view of SS433. Precessional evolution of the jet-disk system. A&A 669, pp. A149. External Links: Document, 2210.09390, ADS entry Cited by: §1, footnote 2.
+- The fourth Uhuru catalog of X-ray sources.. ApJS 38, pp. 357–412. External Links: Document, ADS entry Cited by: §1.
+- Long time-scale numerical simulations of large supercritical accretion discs. MNRAS 540 (3), pp. 2820–2829. External Links: Document, 2505.08859, ADS entry Cited by: §1.
+- SS 433: a phenomenon imitating a Wolf-Rayet star. A&A 445 (3), pp. 1041–1052. External Links: Document, astro-ph/0512190, ADS entry Cited by: §1.
+- Gaia Data Release 3. Summary of the content and survey properties. A&A 674, pp. A1. External Links: Document, 2208.00211, ADS entry Cited by: footnote 1.
+- Continuum observations of the SNR W50 and G 74.9+1.2 at 2695 MHz.. A&A 84, pp. 237–244. External Links: ADS entry Cited by: §1.
+- When microquasar jets and supernova collide: hydrodynamically simulating the SS 433-W 50 interaction.
+MNRAS 414 (4), pp.
+2838–2859.
+External Links: Document, 1101.3486, ADS entry Cited by: §1, §1, §2, §4, §6, §6.
+- Probing the history of SS 433’s jet kinematics via decade-resolution radio observations of W 50. MNRAS 414 (4), pp. 2828–2837. External Links: Document, 1103.5658, ADS entry Cited by: §1.
+- Photometric Mass Estimate for the Compact Component of SS 433: And Yet It Is a Neutron Star. PZ 31 (5), pp. 5. External Links: Document, 1110.5304, ADS entry Cited by: §1.
+- A colour scheme for the display of astronomical intensity images. BASI 39, pp. 289–295. External Links: Document, 1108.5083, ADS entry Cited by: Data availability.
+- An updated catalogue of 310 Galactic supernova remnants and their statistical properties. JApA 46 (1), pp. 14. External Links: Document, 2411.03367, ADS entry Cited by: Figure 1.
+- A 3D Dust Map Based on Gaia, Pan-STARRS 1, and 2MASS. ApJ 887 (1), pp. 93. External Links: Document, 1905.02734, ADS entry Cited by: Figure 9, Appendix C, §2.
+- The GB6 Catalog of Radio Sources. ApJS 103, pp. 427. External Links: Document, ADS entry Cited by: Figure 1, §3, Figure 2.
+- The central X-ray source in SS 433.. ApJ 277, pp. 286–295. External Links: Document, ADS entry Cited by: §1.
+- Acceleration and transport of relativistic electrons in the jets of the microquasar SS 433. Sci 383 (6681), pp. 402–406. External Links: Document, 2401.16019, ADS entry Cited by: §1, §5.2.4.
+- Cosmic rays from microquasars: A narrow component to the CR spectrum?. A&A 390, pp. 751–766. External Links: Document, astro-ph/0204183, ADS entry Cited by: §1.
+- An analysis of the proper motions of SS 433 radio jets.. ApJL 246, pp. L141–L145. External Links: Document, ADS entry Cited by: §1.
+- Structure, strength and polarization changes in radio source SS 433.. Nature 290, pp. 100–107. External Links: Document, ADS entry Cited by: §1.
+- 178 MHz observations of galactic radio sources.. MNRAS 143, pp. 407–435. External Links: Document, ADS entry Cited by: §1.
+- Super-Eddington Accretion Disks around Supermassive Black Holes. ApJ 880 (2), pp. 67. External Links: Document, 1709.02845, ADS entry Cited by: §1.
+- Ultraluminous X-Ray Sources. ARA&A 55 (1), pp. 303–341. External Links: Document, 1703.10728, ADS entry Cited by: §1.
+- X-Ray Polarization of the Eastern Lobe of SS 433. ApJL 961 (1), pp. L12. External Links: Document, 2311.16313, ADS entry Cited by: §1, §1, §3, §5.2.4, §6.
+- Nodding motions of accretion rings and disks : a short-term period inSS 433.. ApJ 260, pp. 780–793. External Links: Document, ADS entry Cited by: §1.
+- Acceleration, radiation and precession in SS 433.. ApJL 236, pp. L127–L130. External Links: Document, ADS entry Cited by: §1.
+- GINGA observations of the X-ray eclipse of SS 433.. PASJ 41, pp. 491–507. External Links: ADS entry Cited by: §1, §2, §5.1.
+- SRG/eROSITA discovery of a radio-faint X-ray candidate supernova remnant SRGe J003602.3+605421 = G121.1-1.9.
+MNRAS 521 (4), pp.
+5536–5556.
+External Links: Document, 2207.00064, ADS entry Cited by: §2.
+- X-ray line formation in the spectrum of SS 433. AstL 38 (7), pp. 443–467. External Links: Document, 1206.0451, ADS entry Cited by: Appendix A.
+- Identification of X-ray lines in the spectrum of the arcsec-scale precessing jets of SS 433. AstL 43 (6), pp. 388–399. External Links: Document, 1701.05884, ADS entry Cited by: §1, §4.
+- Constraints on the Collimated X-ray Emission of SS 433 from the Reflection on Molecular Clouds. AstL 45 (5), pp. 282–298. External Links: Document, 1905.09667, ADS entry Cited by: §1, §5.2.4.
+- Thermal X-ray emission from a baryonic jet: a self-consistent multicolour spectral model.
+MNRAS 455 (2), pp.
+1414–1427.
+External Links: Document, 1510.05563, ADS entry Cited by: Appendix A, §1, Figure 6, §5.1.
+- Is SS 433 a misaligned ultraluminous X-ray source?
+Constraints from its reflected signal in the Galactic plane.
+MNRAS 457 (4), pp.
+3963–3974.
+External Links: Document, 1512.02616, ADS entry Cited by: §1, §5.2.4, footnote 2.
+- SRG/eROSITA view of X-ray reflection in the Central Molecular Zone: a snapshot in September-October 2019. MNRAS 509 (4), pp. 6068–6076. External Links: Document, 2111.07899, ADS entry Cited by: §2.
+- Study of X-ray emission from the S147 nebula by SRG/eROSITA: Supernova-in-the-cavity scenario. A&A 689, pp. A278. External Links: Document, 2401.17261, ADS entry Cited by: §2.
+- Deciphering the Origin of the GeV-TeV Gamma-Ray Emission from SS 433. ApJ 904 (2), pp. 188. External Links: Document, 2008.04515, ADS entry Cited by: §1.
+- Spectrophotometry of W 50 : the SNR around SS 433.. ApJL 242, pp. L77–L81. External Links: Document, ADS entry Cited by: §1.
+- W 50 - a stellar-wind bubble in a three-phase intestellar medium ?. MNRAS 205, pp. 471–485. External Links: Document, ADS entry Cited by: §6.
+- Iron-Line Diagnostics of the Jets of SS 433. PASJ 48, pp. 619–629. External Links: Document, ADS entry Cited by: §1, §5.1.
+- Line Emission Mapper (LEM): Probing the physics of cosmic ecosystems. arXiv, pp. arXiv:2211.09827. External Links: Document, 2211.09827, ADS entry Cited by: §6.
+- Suzaku X-Ray and Optical Spectroscopic Observations of SS 433 in the 2006 April Multiwavelength Campaign. PASJ 62, pp. 323. External Links: Document, 1001.2989, ADS entry Cited by: §1.
+- A geometrical model for the SS 433 system.. MNRAS 210, pp. 279–301. External Links: Document, ADS entry Cited by: §1.
+- Ultrahigh-Energy Gamma-ray Emission Associated with Black Hole-Jet Systems. National Science Review 12 (12), pp. nwaf496. External Links: Document, 2410.08988, ADS entry Cited by: §1.
+- Determining the Nature of the SS 433 Binary from an X-Ray Spectrum during Eclipse. ApJ 650 (1), pp. 338–349. External Links: Document, astro-ph/0605574, ADS entry Cited by: §1, §2, §5.1.
+- The bizarre spectrum of SS 433.. ApJL 230, pp. L41–L45. External Links: Document, ADS entry Cited by: §1.
+- Observations of SS 433. ARA&A 22, pp. 507–536. External Links: Document, ADS entry Cited by: §1.
+- Multiwavelength Observations of the SS 433 Jets. ApJ 775 (1), pp. 75. External Links: Document, 1307.8427, ADS entry Cited by: §1, §1, §2, §5.1.
+- The High-Resolution X-Ray Spectrum of SS 433 Using the Chandra HETGS. ApJ 564 (2), pp. 941–952. External Links: Document, astro-ph/0108206, ADS entry Cited by: Appendix A, §1, §4, Figure 6, §5.1.
+- Evidence of supercritical disc funnel radiation in X-ray spectra of SS 433.
+MNRAS 402 (1), pp.
+479–491.
+External Links: Document, 0910.3186, ADS entry Cited by: Appendix A, §1, Figure 6, §5.1, footnote 2.
+- An Upper Limit on Nickel Overabundance in the Supercritical Accretion Disk Wind of SS 433 from X-ray Spectroscopy.
+AstL 44 (6), pp.
+390–410.
+External Links: Document, 1804.01828, ADS entry Cited by: Appendix A, §1.
+- Diagnostics of Parameters for the X-ray Jets of SS 433 from High-Resolution Chandra Spectroscopy.
+AstL 45 (5), pp.
+299–320.
+External Links: Document, 2005.12416, ADS entry Cited by: Appendix A, §1, §2, Figure 6, §5.1.
+- The X-ray Jets of SS 433 in the Period of Flaring Activity in the Summer of 2018. AstL 48 (7), pp. 389–405. External Links: Document, ADS entry Cited by: §1.
+- NuSTAR reveals the hidden nature of SS433. MNRAS 506 (1), pp. 1045–1058. External Links: Document, 1810.10518, ADS entry Cited by: §1, footnote 2.
+- Rapid variability of the arcsec-scale X-ray jets of SS 433. MNRAS 358 (3), pp. 860–868. External Links: Document, astro-ph/0501097, ADS entry Cited by: §1.
+- Coupled Radio and X-Ray Emission and Evidence for Discrete Ejecta in the Jets of SS 433. ApJ 682 (2), pp. 1141–1151. External Links: Document, 0804.1337, ADS entry Cited by: §1.
+- The SS433 jet from subparsec to parsec scales. A&A 574, pp. A143. External Links: Document, ADS entry Cited by: §6, §6.
+- X-Ray Lobes of W50/SS433 System. AdSpR 25 (3-4), pp. 709–712. External Links: Document, ADS entry Cited by: §1, §3.
+- Continuous Jets and Backflow Models for the Formation of W50/SS 433 in Magnetohydrodynamics Simulations. ApJ 910 (2), pp. 149. External Links: Document, 2102.06728, ADS entry Cited by: §6.
+- Global Structure of Three Distinct Accretion Flows and Outflows around Black Holes from Two-dimensional Radiation-magnetohydrodynamic Simulations.
+ApJ 736 (1), pp.
+2.
+External Links: Document, 1105.5474, ADS entry Cited by: §1.
+- Notes on non-thermal X-ray radiation of radio supernova remnant W50 and collimated radiation of SS433. arXiv, pp. arXiv:1306.4486. External Links: Document, 1306.4486, ADS entry Cited by: §5.2.4.
+- Simulation of kinematics of SS 433 radio jets that interact with the ambient medium. A&A 562, pp. A130. External Links: Document, 1306.3599, ADS entry Cited by: §1, §6.
+- Jets of SS 433 on scales of dozens of parsecs. A&A 599, pp. A77. External Links: Document, 1607.02043, ADS entry Cited by: §1, §6, §6.
+- Particle acceleration and multi-messenger radiation from ultra-luminous X-ray sources: A new class of Galactic PeVatrons.
+A&A 698, pp.
+A188.
+External Links: Document, 2411.08762, ADS entry Cited by: §1.
+- Hydrodynamic Collimation of Precessing Jets. ApJ 417, pp. 170. External Links: Document, ADS entry Cited by: §6, §6.
+- Supercritically accreting stellar mass black holes as ultraluminous X-ray sources. MNRAS 377 (3), pp. 1187–1194. External Links: Document, astro-ph/0609274, ADS entry Cited by: §1, footnote 2.
+- The eROSITA X-ray telescope on SRG. A&A 647, pp. A1. External Links: Document, 2010.03477, ADS entry Cited by: §1, §2, §2.
+- Detection of large-scale X-ray bubbles in the Milky Way halo. Nature 588 (7837), pp. 227–231. External Links: Document, 2012.05840, ADS entry Cited by: §2.
+- Kinematics of the western filament of W50. MNRAS 506 (3), pp. 4263–4274. External Links: Document, ADS entry Cited by: §1.
+- Hard X-Ray Emission from the Eastern Jet of SS 433 Powering the W50 “Manatee” Nebula: Evidence for Particle Reacceleration.
+ApJ 935 (2), pp.
+163.
+External Links: Document, 2207.00573, ADS entry Cited by: §1, §1, §1, §2, §4, §5.2.4, §6.
+- ROSAT and ASCA Observations of W50 Associated with the Peculiar Source SS 433. ApJ 483 (2), pp. 868–881. External Links: Document, ADS entry Cited by: §1, §3.
+- Rossi X-Ray Timing Explorer Observations of the Eastern Lobe of W50 Associated with SS 433. ApJ 512 (2), pp. 784–792. External Links: Document, ADS entry Cited by: §1, §3.
+- Arcsecond-scale X-ray imaging and spectroscopy of SS 433 with the Chandra High-Energy Transmission Grating. PASJ 77 (5), pp. 1113–1125. External Links: Document, 2507.19042, ADS entry Cited by: §1.
+- Energy estimation of high-energy particles associated with the SS 433/W 50 system through radio observation at 1.4 GHz.
+PASJ 73 (3), pp.
+530–544.
+External Links: Document, 2103.05578, ADS entry Cited by: §1.
+- Disk-driven precession in SS 433. ApJL 238, pp. L129–L132. External Links: Document, ADS entry Cited by: §1.
+- X-ray sources in the Aquila-Serpens-Scutum region.. MNRAS 175, pp. 39P–46. External Links: Document, ADS entry Cited by: §1.
+- Diffuse X-ray emission from the jets of SS433. Nature 287 (5785), pp. 806–808. External Links: Document, ADS entry Cited by: §1, §2.
+- Black holes in binary systems. Observational appearance.. A&A 24, pp. 337–355. External Links: ADS entry Cited by: §1.
+- A Catalogue of H II Regions.. ApJS 4, pp. 257. External Links: Document, ADS entry Cited by: Figure 1.
+- Mass loss by SS 433, and its effect on the X-ray and radio emission.. SvA 25, pp. 315–319. External Links: ADS entry Cited by: §1.
+- The optical spectrum of W 50.. PASP 92, pp. 259–261. External Links: Document, ADS entry Cited by: §1.
+- A radio jet in SS433. Nature 282 (5738), pp. 483–484. External Links: Document, ADS entry Cited by: §1.
+- New H-alpha emission stars in the Milky Way.. ApJS 33, pp. 459–469. External Links: Document, ADS entry Cited by: §1.
+- Simultaneous observations of the X-ray and optical eclipse of SS 433 and their impplications.. MNRAS 228, pp. 293–303. External Links: Document, ADS entry Cited by: §1.
+- SRG X-ray orbital observatory. Its telescopes and first scientific results. A&A 656, pp. A132. External Links: Document, 2104.13267, ADS entry Cited by: §1, §2, §5.1.
+- X-Ray Imaging and Spectroscopy Mission. PASJ. External Links: Document, ADS entry Cited by: Appendix A, §6.
+- The First Proper Motion Measurement of the Acceleration Regions in the Large-scale Jets of SS 433 Powering the W50 Nebula.
+ApJL 993 (1), pp.
+L24.
+External Links: Document, 2510.06431, ADS entry Cited by: §1, §3.
+- An early-type binary model for SS 433.. A&A 81, pp. L7–L10. External Links: ADS entry Cited by: §1.
+- Forming short-period Wolf-Rayet X-ray binaries and double black holes through stable mass transfer. MNRAS 471 (4), pp. 4256–4264. External Links: Document, 1701.02355, ADS entry Cited by: §1.
+- SS 433, X-ray binaries and stellar evolution. VA 25 (1), pp. 95–108. External Links: Document, ADS entry Cited by: §1, §1.
+- A numerical simulation of the W 50-SS 433 system. A&A 362, pp. 780–785. External Links: ADS entry Cited by: §1, §1, §6.
+- Galactic Superaccreting X-Ray Binaries as Super-PeVatron Accelerators. ApJL 989 (2), pp. L25. External Links: Document, 2507.21048, ADS entry Cited by: §1.
+- Doppler-shifted X-ray line emission from SS 433.. MNRAS 222, pp. 261–271. External Links: Document, ADS entry Cited by: §1.
+- The X-ray lobes of SS 433.. ApJ 273, pp. 688–696. External Links: Document, ADS entry Cited by: §1, §2, §2, §3.
+- A survey of the continuous radiation from the Galactic System at a frequency of 1390 Mc/s. BAN 14, pp. 215. External Links: ADS entry Cited by: §1.
+- The slaved disc model for SS 433.. MNRAS 193, pp. 707–712. External Links: Document, ADS entry Cited by: §1.
+- On the Absorption of X-Rays in the Interstellar Medium. ApJ 542 (2), pp. 914–924. External Links: Document, astro-ph/0008425, ADS entry Cited by: Appendix B.
+- A Non-Thermal X-Ray Spectrum from the Supernova Remnant W 50. PASJ 46, pp. L109–L113. External Links: ADS entry Cited by: §1.
+- Large-scale outflow structure and radiation properties of super-Eddington flow: Dependence on the accretion rates.
+PASJ 74 (6), pp.
+1378–1395.
+External Links: Document, 2209.01427, ADS entry Cited by: §1.
+- Three-dimensional hydrodynamical simulations of the large-scale structure of W50-SS433. MNRAS 387 (2), pp. 839–844. External Links: Document, ADS entry Cited by: §1, §1, §6.
+- The interaction between the relativistic jets of SS433 and the interstellar medium. MNRAS 192, pp. 731–743. External Links: Document, ADS entry Cited by: §1, §6.
 Appendix A Spectrum of the central source
@@
 2002).
-Some residual emission can also be observed in the iron K complex region between 6 and 7 keV, which could come from the reflection of central source emission on the walls of the optically thick accretion disk funnel (Medvedev & Fabrika 2010) or in the optically thin accretion disk wind (Medvedev et al.
+Some residual emission can also be observed in the iron K complex region between 6 and 7 keV, which could come from the reflection of central source emission on the walls of the optically thick accretion disk funnel (Medvedev and Fabrika 2010) or in the optically thin accretion disk wind (Medvedev et al.
 2018, 2019).
@@
 High-resolution spectroscopy of this emission with the XRISM observatory (Tashiro et al.
-2025) will open new possibilities for characterizing this emission and potentially give access to more intricate diagnostics based on velocity structure and resonant scattering in the jets (Khabibullin & Sazonov 2012), as well as possible signatures of multiphase cooling and ionization state dynamics in supersonically expanding jets prone to thermal instability and fragmentation (Brinkmann et al.
-1988; Brinkmann & Kawai 2000; Khabibullin et al.
+2025) will open new possibilities for characterizing this emission and potentially give access to more intricate diagnostics based on velocity structure and resonant scattering in the jets (Khabibullin and Sazonov 2012), as well as possible signatures of multiphase cooling and ionization state dynamics in supersonically expanding jets prone to thermal instability and fragmentation (Brinkmann et al.
+1988; Brinkmann and Kawai 2000; Khabibullin et al.
 2016).